### v1

# Diophantine Graphs

Abstract
This manuscript introduces Diophantine labeling, a new way of labeling of the vertices for finite simple undirected graphs with some divisibility condition on the edges. Maximal graphs admitting Diophantine labeling are investigated and their number of edges are computed. Some number-theoretic techniques are used to characterize vertices of maximum degree and nonadjacent vertices. Some necessary and sufficient conditions for vertices of equal degrees are found.
Keywords: Graph labeling, Prime graph, Diophantine graph, $p$-adic valuation.
MSC code: 05A10, 05C07, 05C78, 11A05, 11A25, 11B75, 11D04, 11D88.
1 Introduction
A graph labeling is an assignment of real values to the vertices or edges that meet some conditions. The definition of prime labeling was originated by R. Entringer and introduced formally in a paper by R. Tout [12]. A graph $G=(V,E)$ with $n$ vertices has a prime labeling if there exists a bijective map $f:V\rightarrow\{1,2,\dots,n\}$ such that for every $uv\in E$, $(f(u),f(v))=1$. Many results of prime labeling and other types of prime labeling can be found in the dynamic survey of graph labeling by J.A. Gallian [6]. Other variants of prime labeling have been introduced by many researchers (see for example [4], [5], [8], [13]).
This paper gives a generalization for the concept of prime graphs. More precisely, we define a graph $G$ with $n$ vertices to be a Diophantine graph if there exists a bijective map $f:V\rightarrow\{1,2,\dots,n\}$ such that for every $uv\in E$, $(f(u),f(v))\mid n$. Clearly, any prime graph is a Diophantine graph. Moreover, any Diophantine graph with a prime number of vertices is a prime graph. In the maximal Diophantine graph with $n$ vertices, an explicit formulas of the number of edges and the degree of a vertex are proved, the full degree vertices are determined. Furthermore, we provide sufficient and necessary conditions that determine the equality of degrees among non-full degree vertices in the maximal Diophantine graphs.
The paper is organized as follows. Section 2 includes some definitions, preliminaries, lemmas and direct observations. Section 3 presents the number of edges formula in the maximal Diophantine graph with $n$ vertices. Section 4 is partitioned into three subsections, each unveiling important results related to the maximal Diophantine graph with order $n$. Subsection 4.1 covers a characterization of labels of vertices with full degree and a formula that compute the degrees of any vertex. Subsection 4.2 and 4.3 encompass some necessary and sufficient conditions for equality of degrees.
For definitions and terminologies of graph theory, we follow F. Harary [7] and A. Bickle [2]. Also, we follow D. Burton [3] and K.H. Rosen [9] for basic definitions and notations in number theory.
2 Definitions and Preliminaries
In the beginning of this section, we give the definitions of the prime graphs and Diophantine graphs. It also explores the relation between them and some direct observations.
Definition 2.1.
[11], [12] Let $G$ be a graph with $n$ vertices. The graph $G$ is called a prime graph if there exists a bijective map $f:V\rightarrow\{1,2,\dots,n\}$ such that for all $uv\in E$, $(f(u),f(v))=1$. Such a map $f$ is called a prime labeling of $G$. The maximal prime graph with $n$ vertices, denoted by $R_{n}$, is a prime graph such that adding any new edge yields a non-prime graph.
Definition 2.2.
Let $G$ be a graph with $n$ vertices. The graph $G$ is called a Diophantine graph if there exists a bijective map $f:V\rightarrow\{1,2,\dots,n\}$ such that for all $uv\in E$, $(f(u),f(v))\mid n$. Such a map $f$ is called a Diophantine labeling of $G$. The maximal Diophantine graph with $n$ vertices, denoted by $D_{n}$, is a Diophantine graph such that adding any new edge yields a non-Diophantine graph.
Example 2.1.
The Petersen graph $P$ is a Diophantine graph, but the complete graph $K_{5}$ is not.
66662222333344441010101011117777888899995555
11112222333344445555
Figure 1: $P$ is a Diophantine graph while $K_{5}$ is not
Example 2.2.
The following are some maximal Diophantine graphs.
1111222233334444
11112222333344445555
111122223333444455556666
1111555577773333222266664444
11112222333344445555666677778888
Figure 2: Maximal Diophantine Graphs $D_{4},\dots,D_{8}$
It is easy to see that every prime graph is a Diophantine graph, but the converse is not necessarily true, e.g., $D_{4},D_{6},D_{8}$. However, in the case that $n$ is a prime number, a graph is Diophantine if and only if it is a prime graph. Moreover, $D_{n}\cong R_{n}.$ The proof is straightforward using the following lemma.
Lemma 2.1.
Let $n>1$. Then we have $n$ is a prime number  if and only if  for all $a,b\leq n$, $a\neq b$, [$(a,b)=1$ if and only if $(a,b)\mid n$].
Proof.
Let $n$ be a prime number, $a,b<n$ and $a\neq b$. If $(a,b)=1,$ then $(a,b)\mid n$. On the other hand, if $(a,b)\mid n,$ then either $(a,b)=1$ or $(a,b)=n$. In the case that $(a,b)=1$, we have nothing to proof. In the other case that $(a,b)=n,$ we get $a=b=n$ which is a contradiction with the hypothesis.
Conversely, Assume $a\neq b$ and [$(a,b)=1$ if and only if $(a,b)\mid n$] for all $a,b<n$. Suppose, by the way of contradiction, $n>1$ is a composite number. Then there exists prime $p\mid n$ such that $p\leq\frac{n}{2}$. Let $a=p\leq n$, $b=2p\leq n$. Hence, $(a,b)=p\mid n$ and $(a,b)\neq 1,$ which contradicts the assumption. Hence, the result follows. ∎
In the subsequent sections, in any graph $G$ of order $n$, we call $f^{*}(u):=\frac{f(u)}{(f(u),n)}$ the reduced label of a vertex $u$ and we denote the set of all multiples of a positive integer $a$ up to $n$ by $M_{a}$, i.e. $M_{a}:=\{ta\ :\ t=1,2,\dots,\lfloor\frac{n}{a}\rfloor\}.$ Now, we present a crucial definition and an important number-theoretic lemmas that will allow us to formulate and prove our results smoothly.
Definition 2.3.
For a given positive integer $n$ and prime number $p$, the successor of the $p$-adic valuation is defined by $\acute{v}_{p}(n):=v_{p}(n)+1,$ where $v_{p}(n)$ is the $p$-adic valuation of an integer $n.$ The number $p^{\acute{v}_{p}(n)}$ is called the critical prime power number with respect to $p,n$.
Lemma 2.2.
Let $u,v\in V(D_{n})$ and $f$ be a Diophantine labeling of $D_{n}$. Then we have
  • (i)
    for every prime number $p$, $p\mid f^{*}(u)$ if and only if $p^{\acute{v}_{p}(n)}\mid f(u);$
  • (ii)
    if $f(u)\mid f(v)$, then $f^{*}(u)\mid f^{*}(v).$
Proof.
The proof of part $(i)$ is straightforward. Let us prove the second part. The $p$-adic valuation function [1] is extended normally to rational numbers as follows
$v_{p}\left(\frac{b}{a}\right):=v_{p}(b)-v_{p}(a),$
for any $a,b\in\mathbb{Z}^{+}$ and for every prime number $p.$ Let $u,v\in V(D_{n})$ such that $f(u)\mid f(v).$ Using the basic properties of the $p$-adic valuation function, we have
$\begin{split}v_{p}\left(\frac{f(v)}{f(u)}\frac{(f(u),n)}{(f(v),n)}\right)&=v_{p}(f(v))-v_{p}(f(u))+v_{p}\big{(}(f(u),n)\big{)}-v_{p}\big{(}(f(v),n)\big{)}.\\ &=v_{p}(f(v))-v_{p}(f(u))+\min\big{\{}v_{p}(f(u)),v_{p}(n)\big{\}}-\min\big{\{}v_{p}(f(v)),v_{p}(n)\big{\}}.\end{split}$
Therefore, one can easily check that
$v_{p}\left(\frac{f(v)}{f(u)}\frac{(f(u),n)}{(f(v),n)}\right)\geq 0$
in all possible cases between $v_{p}(f(u)),v_{p}(n)$ and $v_{p}(f(v))$, namely
$v_{p}(n)<v_{p}(f(u)),\quad v_{p}(f(u))\leq v_{p}(n)\leq v_{p}(f(v))\quad\mbox{and}\quad v_{p}(f(v))<v_{p}(n).$
Hence,
$\frac{f(u)}{(f(u),n)}\ \Bigg{|}\ \frac{f(v)}{(f(v),n)},$
which completes the proof. ∎
In the proof of Lemma 2.3, we will use the following identity: $\frac{1}{j}-\frac{1}{j+1}=\left(1+\frac{2}{j}\right)\left[\frac{1}{j+1}-\frac{1}{j+2}\right]$, $j\in\mathbb{Z}^{+}$.
Lemma 2.3.
For every $n\in\mathbb{Z}^{+}$, there exists $j\in\left\{\lfloor\sqrt{n}\rfloor,\lfloor\sqrt{n}\rfloor+1,\ldots,n\right\}$ such that
$\frac{n}{j+1}<\lfloor\sqrt{n}\rfloor\leq\frac{n}{j}\quad\mbox{and}\quad\frac{n}{j}-\frac{n}{j+1}<1.$
Proof.
Clearly, we have
$\begin{cases}\lfloor\sqrt{n}\rfloor\in\left(\frac{n}{\lfloor\sqrt{n}\rfloor+1},\frac{n}{\lfloor\sqrt{n}\rfloor}\right]\quad\mbox{and}\quad\frac{n}{\lfloor\sqrt{n}\rfloor}-\frac{n}{\lfloor\sqrt{n}\rfloor+1}<1,&\quad\mbox{if}\quad n<\lfloor\sqrt{n}\rfloor(\lfloor\sqrt{n}\rfloor+1);\\ &\\ \lfloor\sqrt{n}\rfloor\notin\left(\frac{n}{\lfloor\sqrt{n}\rfloor+1},\frac{n}{\lfloor\sqrt{n}\rfloor}\right]\quad\mbox{and}\quad\frac{n}{\lfloor\sqrt{n}\rfloor}-\frac{n}{\lfloor\sqrt{n}\rfloor+1}\geq 1,&\quad\mbox{if}\quad n\geq\lfloor\sqrt{n}\rfloor(\lfloor\sqrt{n}\rfloor+1).\\ \end{cases}$
Since $\lfloor\sqrt{n}\rfloor<\frac{n}{\lfloor\sqrt{n}\rfloor}$ and $\frac{n}{j}-\frac{n}{j+1}>\frac{n}{j+1}-\frac{n}{j+2},$ therefore there exists $j\geq\lfloor\sqrt{n}\rfloor$ such that
$\lfloor\sqrt{n}\rfloor\in\left(\frac{n}{j+2},\frac{n}{j+1}\right]\quad\mbox{and}\quad\frac{n}{j+1}-\frac{n}{j+2}\leq\frac{n}{\lfloor\sqrt{n}\rfloor+1}-\frac{n}{\lfloor\sqrt{n}\rfloor+2}<1,$
which complete the proof. ∎
Lemma 2.4.
Let $a,b$ and $n\in\mathbb{Z}^{+}$. If $\lfloor\frac{n}{a}\rfloor=\lfloor\frac{n}{b}\rfloor$ and $a\neq b$, then $ab>n$.
Proof.
Let $a,b$ and $n\in\mathbb{N}$ such that $\lfloor\frac{n}{a}\rfloor=\lfloor\frac{n}{b}\rfloor$. Therefore, there exists $j\in\{1,2,\dots,n\}$ such that $\frac{n}{j+1}<a,b\leq\frac{n}{j}$. Consequently, $\frac{n}{j}-\frac{n}{j+1}>1$, since $a\neq b$. Using Lemma 2.3, we get that there exists $i\in\left\{\lfloor\sqrt{n}\rfloor,\lfloor\sqrt{n}\rfloor+1,\dots,n\right\}$ such that
$\frac{n}{i+1}<\lfloor\sqrt{n}\rfloor\leq\frac{n}{i}\quad\mbox{and}\quad\frac{n}{i}-\frac{n}{i+1}<1.$
Thus, $a,b>\lfloor\sqrt{n}\rfloor$. Hence, $ab>n$. ∎
Lemma 2.5.
Let $a,b,n\in\mathbb{Z}^{+}$ and let $p_{1},p_{2},q_{1},q_{2}$ be distinct prime numbers and $\alpha_{1},\alpha_{2},\beta_{1},\beta_{2}$ be positive integers. If $\frac{a}{(a,n)}=p_{1}^{\alpha_{1}}p_{2}^{\alpha_{2}}$ and $\frac{b}{(b,n)}=q_{1}^{\beta_{1}}q_{2}^{\beta_{2}}$ such that
$\frac{n}{i+1}<p_{1}^{\acute{v}_{p_{1}}(n)},q_{1}^{\acute{v}_{q_{1}}(n)}\leq\frac{n}{i}\quad\mbox{and}\quad\frac{n}{j+1}<p_{2}^{\acute{v}_{p_{2}}(n)},q_{2}^{\acute{v}_{q_{2}}(n)}\leq\frac{n}{j}$ (1)
for some $i,j=1,2,\dots,n$, then
$a>n\quad\mbox{or}\quad b>n.$
Proof.
Let $\frac{a}{(a,n)}=p_{1}^{\alpha_{1}}p_{2}^{\alpha_{2}}$ and $\frac{b}{(b,n)}=q_{1}^{\beta_{1}}q_{2}^{\beta_{2}}$ such that $p_{1},p_{2},q_{1},q_{2}$ are distinct prime numbers satisfying equation (1), where $\alpha_{1},\alpha_{2},\beta_{1},\beta_{2}>0$ are positive integers. Consequently,
$\left\lfloor\frac{n}{p_{1}^{\acute{v}_{p_{1}}(n)}}\right\rfloor=\left\lfloor\frac{n}{q_{1}^{\acute{v}_{q_{1}}(n)}}\right\rfloor\quad\mbox{and}\quad\left\lfloor\frac{n}{p_{2}^{\acute{v}_{p_{2}}(n)}}\right\rfloor=\left\lfloor\frac{n}{q_{2}^{\acute{v}_{q_{2}}(n)}}\right\rfloor.$ (2)
Since $\frac{a}{(a,n)}=p_{1}^{\alpha_{1}}p_{2}^{\alpha_{2}}$ and $\frac{b}{(b,n)}=q_{1}^{\beta_{1}}q_{2}^{\beta_{2}},$ therefore Lemmas 2.2 part $(i)$ implies that
$p_{1}^{\acute{v}_{p_{1}}(n)}p_{2}^{\acute{v}_{p_{2}}(n)}\ \Big{|}\ a\quad\mbox{and}\quad q_{1}^{\acute{v}_{q_{1}}(n)}q_{2}^{\acute{v}_{q_{2}}(n)}\ \Big{|}\ b.$ (3)
In the case $i=j,$ equation (1) implies that
$\left\lfloor\frac{n}{p_{1}^{\acute{v}_{p_{1}}(n)}}\right\rfloor=\left\lfloor\frac{n}{q_{1}^{\acute{v}_{q_{1}}(n)}}\right\rfloor=\left\lfloor\frac{n}{p_{2}^{\acute{v}_{p_{2}}(n)}}\right\rfloor=\left\lfloor\frac{n}{q_{2}^{\acute{v}_{q_{2}}(n)}}\right\rfloor.$
Therefore, using Lemma 2.4 and equation (3), we get
$n<p_{1}^{\acute{v}_{p_{1}}(n)}p_{2}^{\acute{v}_{p_{2}}(n)}\leq a\quad\mbox{and}\quad n<q_{1}^{\acute{v}_{q_{1}}(n)}q_{2}^{\acute{v}_{q_{2}}(n)}\leq b.$
If $i>j$, then equation (3) implies that
$p_{1}^{\acute{v}_{p_{1}}(n)}<p_{2}^{\acute{v}_{p_{2}}(n)},q_{2}^{\acute{v}_{q_{2}}(n)}\quad\mbox{and}\quad q_{1}^{\acute{v}_{q_{1}}(n)}<p_{2}^{\acute{v}_{p_{2}}(n)},q_{2}^{\acute{v}_{q_{2}}(n)}.$
Therefore, using Lemma 2.4 and equations (2) and (3), we get
$n<p_{1}^{\acute{v}_{p_{1}}(n)}q_{1}^{\acute{v}_{q_{1}}(n)}<p_{1}^{\acute{v}_{p_{1}}(n)}p_{2}^{\acute{v}_{p_{2}}(n)}\leq a.$
In a very similar way, one can prove that, if $i<j$, then
$n<p_{2}^{\acute{v}_{p_{2}}(n)}q_{2}^{\acute{v}_{q_{2}}(n)}<q_{1}^{\acute{v}_{q_{1}}(n)}q_{2}^{\acute{v}_{q_{2}}(n)}\leq b.$
Hence, the result follows. ∎
Lemma 2.6.
Let $a,b,t,n\in\mathbb{Z}^{+}$. If $\left\lfloor\frac{n}{a}\right\rfloor=\left\lfloor\frac{n}{b}\right\rfloor$, then $\left\lfloor\frac{n}{ta}\right\rfloor=\left\lfloor\frac{n}{tb}\right\rfloor$.
The proof of Lemma 2.6 is a direct application of the fact $\left\lfloor\frac{x}{n}\right\rfloor=\left\lfloor\frac{\lfloor x\rfloor}{n}\right\rfloor$ for any $x\in\mathbb{R}$ and $n\in\mathbb{Z}^{+}.$ Lemmas 2.5 and 2.6 will be used in the proof of Theorem 4.5.
3 Number of Edges of $D_{n}$
The following lemma is a characterization of nonadjacent vertices in the maximal Diophantine graphs $D_{n}$.
Lemma 3.1.
Suppose $D_{n}$ is the maximal Diophantine graph of order $n$ with a labeling $f$. For each $u,v\in V(D_{n}),$ $uv\notin E(D_{n})$ if and only if there exists a prime number $p$ such that
$f(u),f(v)\in M_{p^{\acute{v}_{p}(n)}}:=\left\{kp^{\acute{v}_{p}(n)}:\ k=1,2,\dots,\left\lfloor\frac{n}{p^{\acute{v}_{p}(n)}}\right\rfloor\right\}.$
Proof.
Suppose $D_{n}$ is the maximal Diophantine graph of order $n$ with a labeling $f$ and let $u,v\in V(D_{n})$. Then we have
$\begin{split}uv\notin E(D_{n})&\quad\mbox{if and only if}\quad(f(u),f(v))\nmid n\\ &\quad\mbox{if and only if}\quad v_{p}\left(\big{(}f(u),f(v)\big{)}\right)>v_{p}(n)\quad\mbox{for some prime number $p$}\\ &\quad\mbox{if and only if}\quad\min\left\{v_{p}(f(u)),\ v_{p}(f(v))\right\}\geq v_{p}(n)+1=\acute{v}_{p}(n)\quad\mbox{for some prime number $p$}\\ &\quad\mbox{if and only if}\quad v_{p}(f(u))\geq\acute{v}_{p}(n)\quad\mbox{and}\quad v_{p}(f(v))\geq\acute{v}_{p}(n)\quad\mbox{for some prime number $p$}\\ &\quad\mbox{if and only if}\quad p^{\acute{v}_{p}(n)}\mid f(u)\quad\mbox{and}\quad p^{\acute{v}_{p}(n)}\mid f(v)\quad\mbox{for some prime number $p$}\\ &\quad\mbox{if and only if}\quad f(u),\ f(v)\in M_{p^{\acute{v}_{p}(n)}}\quad\mbox{for some prime number $p$.}\hskip 139.41832pt\square\end{split}$
It is clear that from the Lemma 3.1, if such a prime number $p$ exists, then $p<\frac{n}{2}$. Any edge in $D_{n}$ represents a solvable linear Diophantine equation $ax+by=n$ and any non-edge in $D_{n}$ represents an unsolvable linear Diophantine equation $ax+by=n$, more precisely, using lemma 3.1, we have for each $u,v\in V(D_{n})$, $uv\notin E(D_{n})$ if and only if the linear equation $f(u)x+f(v)y=n$ is unsolvable.
Theorem 3.1.
The number of edges of the maximal Diophantine graph $D_{n}$ with order $n$ is given by
$C(n,2)-\sum_{1\leq i\leq\pi(n)}C\Bigg{(}\Bigg{\lfloor}\frac{n}{p_{i}^{\acute{v}_{p_{i}}(n)}}\Bigg{\rfloor},2\Bigg{)}+\sum_{1\leq i<j\leq\pi(n)}C\Bigg{(}\Bigg{\lfloor}\frac{n}{p_{i}^{\acute{v}_{p_{i}}}(n)p_{j}^{\acute{v}_{p_{j}}(n)}}\Bigg{\rfloor},2\Bigg{)}-\dots+(-1)^{\pi(n)}C\left(\left\lfloor\frac{n}{\prod\limits_{1\leq i\leq\pi(n)}p_{i}^{\acute{v}_{p_{i}}(n)}}\right\rfloor,2\right),$
where $p_{1},p_{2},\dots,p_{\pi(n)}<n$ are distinct prime numbers, $\pi(n)$ is the number of prime numbers not exceeding $n$ and $C(n,r)$ is the binomial coefficient with $C(n,r):=0$ if $n<r$ and $C(n,r):=\binom{n}{r}$ if $n\geq r$.
Proof. Let $D_{n}$ be the maximal Diophantine graph with $n$ vertices, $f$ be a Diophantine labeling of $D_{n}$ and $p_{i}<n,\ i=1,2,\dots,\pi(n)$ be distinct prime numbers. Using Lemma 3.1, we get for each $u,v\in V(D_{n})$, $uv\notin E(D_{n})$ if and only if $f(u),f(v)\in M_{p_{i}^{\acute{v}_{p_{i}}(n)}}$ for some $i=1,2,\dots,\pi(n)$. Define
$S_{p_{i}}:=\big{\{}\{u,v\}\subset V(D_{n}):f(u),f(v)\in M_{p_{i}^{\acute{v}_{p_{i}}(n)}},\ f(u)\neq f(v)\big{\}},$
$i=1,2,\dots,\pi(n)$. So, we have the following equation
$\begin{split}|E(D_{n})|&=|E(K_{n})|-|E(\overline{D_{n}})|,\ \mbox{where $\overline{D_{n}}$ is the complement of $D_{n}$.}\\ &=|E(K_{n})|-\left|\bigcup\limits_{1\leq i\leq\pi(n)}S_{p_{i}}\right|.\end{split}$
Using the inclusion-exclusion principle (see [10]), we have
$|E(D_{n})|=C(n,2)-\left[\sum_{1\leq i\leq\pi(n)}|S_{p_{i}}|-\sum_{1\leq i<j\leq\pi(n)}|S_{p_{i}}\cap S_{p_{j}}|+\dots+(-1)^{\pi(n)-1}\left|\bigcap\limits_{1\leq i\leq\pi(n)}S_{p_{i}}\right|\right].$
Therefore,
$\begin{split}|E(D_{n})|&=C(n,2)-\sum_{1\leq i\leq\pi(n)}C\left(\left|M_{p_{i}^{\acute{v}_{p_{i}}(n)}}\right|,2\right)+\sum_{1\leq i<j\leq\pi(n)}C\left(\left|M_{p_{i}^{\acute{v}_{p_{i}}(n)}}\cap M_{p_{j}^{\acute{v}_{p_{j}}(n)}}\right|,2\right)-\\ &\qquad\qquad\quad\dots+(-1)^{\pi(n)}C\left(\left|\bigcap\limits_{1\leq i\leq\pi(n)}M_{p_{i}^{\acute{v}_{p_{i}}(n)}}\right|,2\right).\\ &=C(n,2)-\sum_{1\leq i\leq\pi(n)}C\left(\left\lfloor\frac{n}{p_{i}^{\acute{v}_{p_{i}}(n)}}\right\rfloor,2\right)+\sum_{1\leq i<j\leq\pi(n)}C\left(\left\lfloor\frac{n}{p_{i}^{\acute{v}_{p_{i}}(n)}p_{j}^{\acute{v}_{p_{j}}(n)}}\right\rfloor,2\right)-\\ &\qquad\qquad\quad\dots+(-1)^{\pi(n)}C\left(\left\lfloor\frac{n}{\prod\limits_{1\leq i\leq\pi(n)}p_{i}^{\acute{v}_{p_{i}}(n)}}\right\rfloor,2\right).\end{split}$
4 Degrees of Vertices
4.1 The Degree of a Vertex in $D_{n}$
Theorem 4.1.
Suppose $D_{n}$ is the maximal Diophantine graph of order $n$ with a labeling $f$. For each $u\in V(D_{n})$, $p$ is a prime number,
$\deg(u)=n-1\quad\mbox{if and only if}\quad f(u)\mid n\quad\mbox{{or}}\quad\frac{n}{2}<f(u)=p^{\acute{v}_{p}(n)}<n,$
where the exclusive or will be typed in bold while the inclusive or is typed as usual.
Proof.
Assume that $D_{n}$ is the maximal Diophantine graph of order $n$ with a labeling $f$ and $u\in V(D_{n})$ with $\deg(u)=n-1$. By the way of contradiction, suppose
$f(u)\nmid n\quad\mbox{and}\quad\left(f(u)\neq p^{\acute{v}_{p}(n)}\quad\mbox{or}\quad\frac{n}{2}\nless f(u)=p^{\acute{v}_{p}(n)}\nless n\right),$
where $p$ is a prime number. In case of $f(u)=p^{\acute{v}_{p}(n)}$ and $1<f(u)<\frac{n}{2}$, we have
$(f(u),f(v))=(f(u),2f(u))=\left(p^{\acute{v}_{p}(n)},2p^{\acute{v}_{p}(n)}\right)=p^{\acute{v}_{p}(n)}\nmid n,$
where $v\in V(D_{n})$ such that $f(v)=2f(n)<n$, which contradicts the full degree of $u$. In the other case, when $f(u)\neq p^{\acute{v}_{p}(n)}$ and $f(u)\nmid n$, we have that there exists a prime number $p\mid f(u)$ such that $v_{p}(f(u))>v_{p}(n).$ Then we get
$v_{p}\left(f(u)\right)\geq v_{p}(n)+1=\acute{v}_{p}(n),$
so $p^{\acute{v}_{p}(n)}\mid f(u)$. Therefore, one can see that
$\left(f(u),p^{\acute{v}_{p}(n)}\right)=p^{\acute{v}_{p}(n)}\nmid n,$
which is a contradiction of the full degree of $u$. Therefore, one can see that
$f(u)\nmid n\quad\mbox{{or}}\quad f(u)=p^{\acute{v}_{p}(n)},\quad\frac{n}{2}<p^{\acute{v}_{p}(n)}<n.$
Conversely, in the first case of $f(u)\mid n$, we have for each $v\in V(D_{n})$, $(f(u),f(v))\mid n$. In the second case of $f(u)=p^{\acute{v}_{p}(n)}$ and $\frac{n}{2}<f(u)<n$, where $p$ is a prime number, we get any proper multiple of $f(u)$ is bigger than $n$. Since $f(u)=p^{\acute{v}_{p}(n)}$, therefore for all $v\in V(D_{n})$, $(f(u),f(v))\mid p^{v_{p}(n)}$, which means for all $v\in V(D_{n})$, $(f(u),f(v))\mid n$. In the both cases, we get $\deg(u)=n-1$. ∎
Similar results can be deduced in the case of the maximal prime graphs $R_{n}$ for Lemma 3.1 and Theorem 4.1. The properties of the maximal Diophantine graphs $D_{n}$ given in Lemma 3.1 and Theorem 4.1 are essential in what follows.
Corollary 4.1.
There are infinitely many Diophantine graphs which are not prime graphs.
Proof.
Let $D_{p^{k}}$ be the maximal Diophantine graphs with $p^{k}$ vertices, where $k>1$ and $p$ is a prime number. $D_{p^{k}}$ can not be a prime graph because the number full degree vertices in $R_{p^{k}}$ that is equal to
[…]
is less than of the number full degree vertices in $D_{p^{k}}$ which is at least
[…]
Hence, the proof follows. ∎
Lemma 4.1.
Suppose $D_{n}$ is the maximal Diophantine graph of order $n$ with a labeling $f$ and $u\in V(D_{n})$. If $f^{*}(u)=p^{k}$, for some $p$ is a prime number and $k\in\mathbb{N}$, then
$\deg(u)=\left\{\begin{array}[]{ll}n-1,&\hbox{$k=0.$}\\ n-\Big{\lfloor}\frac{n}{p^{\acute{v}_{p}(n)}}\Big{\rfloor},&\hbox{$k>0.$}\end{array}\right.$
Proof.
Let $D_{n}$ be the maximal Diophantine graph of order $n$ with a labeling $f$. Let $u\in V(D_{n})$ such that $f^{*}(u)=p^{k}$, where $p$ is a prime number and $k\geq 0$. In the case of $k=0$, we have $f^{*}(u)=1$. Since $f^{*}(u)=\frac{f(u)}{(f(u),n)}$, we have $f(u)\mid n$. Therefore, using Theorem 4.1, $\deg(u)=n-1$. In the other case of $k>0$, we have
[…]
Let $s\in V(D_{n})-\{u\}$ such that
[…]
Therefore,
$\deg(u)=n-1-\Big{(}\Big{|}M_{p^{\acute{v}_{p}(n)}}\Big{|}-1\Big{)}=n-\left\lfloor\frac{n}{p^{\acute{v}_{p}(n)}}\right\rfloor.$
Hence, we have the proof. ∎
Example 4.1.
In $D_{20}$, let $f(u)=18$ be the label of a vertex $u\in D_{20}$. Then $f^{*}(u)=9=3^{2}$ and we have $k=2>0$. Therefore, Using Lemma 4.1, we get the following
$\deg(u)=20-\left\lfloor\frac{20}{3^{\acute{v}_{3}(20)}}\right\rfloor=14.$
Another way to compute the degree of $u$ with label $18$. The neighborhood of $u$ is the set of adjacent vertices of $u$ that has labels belong to the set $\{1,2,4,5,7,8,10,11,13,14,16,17,19,20\}$. Therefore, $\deg(u)=14$.
In a very similar way of Lemma 4.1, using the inclusion-exclusion principle and the identity $\sum\limits_{i=0}^{r}(-1)^{r}\binom{r}{i}=0$, one can prove the following theorem.
Theorem 4.2.
Suppose $D_{n}$ is the maximal Diophantine graph with order $n$, $f$ is a Diophantine labeling and $u\in V(D_{n})$. If $f^{*}(u)=\prod\limits_{i=1}^{r}p_{i}^{k_{i}}$, where $p_{i}$, $i=1,2,\dots,r$ are distinct prime numbers and $k_{i}\geq 0$, $i=1,2,\dots,r$, then
$\deg(u)=\left\{\begin{array}[]{ll}n-1,&\hbox{$f^{*}(u)=1.$}\\ n-\sum\limits_{1\leq i\leq r}\bigg{\lfloor}\frac{n}{p_{i}^{\acute{v}_{p_{i}}(n)}}\bigg{\rfloor}+\sum\limits_{1\leq i.j\leq r}\bigg{\lfloor}\frac{n}{p_{i}^{\acute{v}_{p_{i}}(n)}p_{j}^{\acute{v}_{p_{j}}(n)}}\bigg{\rfloor}-\dots+(-1)^{r}\Bigg{\lfloor}\frac{n}{\prod\limits_{1\leq i\leq r}p_{i}^{\acute{v}_{p_{i}}(n)}}\Bigg{\rfloor},&\hbox{$f^{*}(u)>1.$}\end{array}\right.$
Similar results of Theorems 3.1, 4.2 can be obtained in the maximal prime graph $R_{n}$, by putting $v_{p}(n)=0$.
4.2 Vertices of Equal Degrees $<n-1$
Lemma 4.2.
Suppose $D_{n}$ is the maximal Diophantine graph of order $n$ with a labeling $f$ and $u,v\in V(D_{n})$. If $f(u)\mid f(v)$, then $N(u)\supseteq N(v)$, where $N(s)$ defines the neighborhood of $s$ as the set of all vertices in $D_{n}$ that join the vertex $s$.
Proof.
The proof is straightforward. ∎
Lemma 4.3.
Suppose $D_{n}$ is the maximal Diophantine graph of order $n$ with a labeling $f$. Let $u,v\in V(D_{n})$ such that $f(u)\mid f(v)$. If $\deg(u)=\deg(v)$, then $N(u)=N(v)$.
Proof.
Suppose $D_{n}$ is the maximal Diophantine graph of order $n$ with a labeling $f$. Let $u,v\in V(D_{n})$ such that
$f(u)\mid f(v)\quad\mbox{and}\quad\deg(u)=\deg(v).$
Let $s\in V(D_{n})-\{u,v\}$ such that $sv\in E(D_{n})$. Then, using Lemma 4.2, $su\in E(D_{n})$. Hence, $N(u)\supseteq N(v)$. On the other hand, let $s\in V(D_{n})-\{u,v\}$ such that $su\in E(D_{n})$. Suppose by contrary that $sv\notin E(D_{n})$. Since $f(u)\mid f(v)$ and using Lemma 4.2, therefore $N(u)\supset N(v)$. Then we have $\deg(u)>\deg(v)$, which contradicts the hypothesis of $\deg(u)=\deg(v)$. Therefore, $sv\in E(D_{n})$. Consequently, $N(u)\subseteq N(v)$. Hence, $N(u)=N(v)$. ∎
It is clear that, if $N(u)=N(v)$, then $\deg(u)=\deg(v)$. Thus, under the assumption that $f(u)\mid f(v)$, we have
$N(u)=N(v)\quad\mbox{ if and only if}\quad\deg(u)=\deg(v).$
Furthermore, the following example shows that the condition $f(u)\mid f(v)$ is necessary for Lemma 4.3.
Example 4.2.
In $D_{20}$, let $f(u)=7$ and $f(v)=8$, where $u,v\in V(D_{20})$. $N(u)=V(D_{20})-\{u_{1},u\}$, where $u_{1}\in V(D_{20})$ such that $f(u_{1})=14$ and $N(v)=V(D_{20})-\{v_{1},v\}$, where $v_{1}\in V(D_{20})$ such that $f(v_{1})=16$. Then we have $\deg(u)=\deg(v)$ while $f(u)\nmid f(v)$, but $N(u)\neq N(v)$.
Lemma 4.4.
Suppose $D_{n}$ is the maximal Diophantine graph of order $n$ with a labeling $f$ and $u\in V(D_{n})$. If $deg(u)<n-1$, then there exists a prime number $p$ such that $p^{\acute{v}_{p}(n)}\mid f(u)$.
Proof.
Suppose $D_{n}$ is the maximal Diophantine graph of order $n$ with a labeling $f$ and let $u\in V(D_{n})$ with $deg(u)<n-1$. Then there exists a vertex $s\in V(D_{n})$ such that $su\notin E(D_{n})$. Therefore, using Lemma 3.1, there exists a prime number $p$ such that $f(s),f(u)\in M_{p^{\acute{v}_{p}(n)}}$. Hence, there exists a prime number $p$ such that $p^{\acute{v}_{p}(n)}\mid f(u)$. ∎
Theorem 4.3.
Suppose $D_{n}$ is the maximal Diophantine graph of order $n$ with a labeling $f$. Let $f(u)=p^{t}$ and $f(v)=q^{k}$ for some $t,k\geq 0$, where $u,v\in V(D_{n})$ with $\deg(u)<n-1$, $\deg(v)<n-1$ and $p,q$ are two prime numbers. Then we have
$\deg(u)=\deg(v)\quad\mbox{if and only if}\quad\acute{v}_{p}(n)\leq t,\ \acute{v}_{q}(n)\leq k\quad\mbox{and}\quad\left\lfloor\frac{n}{p^{\acute{v}_{p}(n)}}\right\rfloor=\left\lfloor\frac{n}{q^{\acute{v}_{q}(n)}}\right\rfloor.$
Proof.
Suppose $D_{n}$ is the maximal Diophantine graph of order $n$ with a labeling $f$. Let $f(u)=p^{t}$ and $f(v)=q^{k}$ for some $t,k>0$, where $p,q$ are primes numbers and $u$ $v\in V(D_{n})$ with $deg(u)=deg(v)$, $deg(u)<n-1$, $deg(v)<n-1$. Therefore, using Lemma 4.4, we have
$p^{\acute{v}_{p}(n)}\mid f(u)=p^{t}\quad\mbox{and}\quad q^{\acute{v}_{q}(n)}\mid f(v)=q^{k}.$
Consequently, $\acute{v}_{p}(n)\leq t$ and $\acute{v}_{q}(n)\leq k$. By using Lemma 4.1, we get
$deg(u)=n-\left\lfloor\frac{n}{p^{\acute{v}_{p}(n)}}\right\rfloor\quad\mbox{and}\quad\deg(v)=n-\left\lfloor\frac{n}{q^{\acute{v}_{q}(n)}}\right\rfloor.$
Hence, we obtain the following equation
$\left\lfloor\frac{n}{p^{\acute{v}_{p}(n)}}\right\rfloor=\left\lfloor\frac{n}{q^{\acute{v}_{q}(n)}}\right\rfloor.$
Conversely, let $\left\lfloor\frac{n}{p^{\acute{v}_{p}(n)}}\right\rfloor=\left\lfloor\frac{n}{q^{\acute{v}_{q}(n)}}\right\rfloor$ and $\acute{v}_{p}(n)\leq t$, $\acute{v}_{q}(n)\leq k$. Then we get $p^{\acute{v}_{p}(n)}\mid f(u)\ \mbox{and}\ q^{\acute{v}_{q}(n)}\mid f(v).$ Thus, using Lemma 4.1, we have the following two equations
$\deg(u)=n-\left\lfloor\frac{n}{p^{\acute{v}_{p}(n)}}\right\rfloor\quad\mbox{and}\quad\deg(v)=n-\left\lfloor\frac{n}{q^{\acute{v}_{q}(n)}}\right\rfloor.$
Hence, the result follows. ∎
Corollary 4.2.
Suppose $D_{n}$ is the maximal Diophantine graph of order $n$ with a labeling $f.$ Let $u,v$ be two vertices with $\deg(u)<n-1$, $\deg(v)<n-1$ such that $f(u)=p^{t}$, $f(v)=p^{k}$ for some $t,k\geq 0$, where $p$ is a prime number. Then we have
$\deg(u)=\deg(v)\quad\mbox{if and only if}\quad\acute{v}_{p}(n)\leq\min\{t,k\}.$
Theorem 4.4.
Suppose $D_{n}$ is the maximal Diophantine graph of order $n$ with a labeling $f.$ Let $u,v$ be two vertices such that $f(v)=tf(u)$ for some $t\geq 1$ and $f(v)$ is not a prime power. Then we have
$\deg(u)=\deg(v)\quad\mbox{if and only if for every prime number}\ p\mid t,\quad\acute{v}_{p}(n)\leq v_{p}(f(u))\quad\mbox{{or}}\quad\acute{v}_{p}(n)>v_{p}((f(v))).$
Proof.
Suppose $D_{n}$ is the maximal Diophantine graph of order $n$ with a labeling $f$. Let $u,v\in V(D_{n})$ such that $f(v)=tf(u)$ for some $t\geq 1$, $f(v)$ is not a prime power. Proof by contrapositive, suppose that there exists a prime number $p_{0}\mid t$ such that
$\acute{v}_{p_{0}}(n)\leq v_{p_{0}}(f(u))\quad\mbox{and}\quad\acute{v}_{p_{0}}(n)>v_{p_{0}}((f(v))),$
or
$\acute{v}_{p_{0}}(n)>v_{p_{0}}(f(u))\quad\mbox{and}\quad\acute{v}_{p_{0}}(n)\leq v_{p_{0}}(f(v)).$
The first case contradicts the assumptions $f(u)\mid f(v)$. Therefore, let us consider the second case, which is
$v_{p_{0}}(f(u))<\acute{v}_{p_{0}}(n)\leq v_{p_{0}}(f(v)).$ (4)
Let a vertex $s\in V(D_{n})$ such that $f(s)=p_{0}^{\acute{v}_{p_{0}}(n)}$. Therfore, using equation (4), $p_{0}^{\acute{v}_{p_{0}}(n)}\nmid p_{0}^{\acute{v}_{p_{0}}(f(u))}$. Since $p_{0}^{\acute{v}_{p_{0}}(f(u))}\mid f(u)$, we get $f(s)=p_{0}^{\acute{v}_{p_{0}}(n)}\nmid f(u)$. Consequently, $s\neq u$. Also, $s\neq v$ for $f(v)$ is not a prime power. Moreover, using equation (4), we get
$(f(u),f(s))\mid n\quad\mbox{while}\quad(f(v),f(s))\nmid n.$
Thus,
$su\in E(D_{n})\quad\mbox{while}\quad sv\notin E(D_{n}).$
Since $f(u)\mid f(v)$ and using Lemma 4.2, we have $N(u)\supset N(v)$. Hence, $\deg(u)\neq\deg(v)$.
Conversely, suppose that, by contrapositive, $deg(u)\neq deg(v)$. Then, using Lemma 4.2, $N(u)\supset N(v)$. So, there exists a vertex $s\in V(D_{n})-\{u,v\}$ such that
$s\in N(u)\quad\mbox{and}\quad s\notin N(v).$
Consequently,
$(f(u),f(s))\mid n\quad\mbox{and}\quad(f(v),f(s))\nmid n.$ (5)
Then, using Lemma 3.1, there exists a prime number $p_{0}$ such that
$f(v),f(s)\in M_{p_{0}^{\acute{v}_{p_{0}}(n)}}\quad\mbox{and}\quad p_{0}^{\acute{v}_{p_{0}}(n)}\nmid f(u).$ (6)
Let $f(s)=p_{0}^{\acute{v}_{p_{0}}(n)}$. Then, using equation 6, we have $s\neq u$. Since $f(v)=tf(u)$, we get $p_{0}\mid t$ and $s\neq v$. Therefore, using equation 5, we have the following
$v_{p_{0}}\left((f(u),f(s))\right)\leq v_{p_{0}}(n)\quad\mbox{and}\quad v_{p_{0}}\left((f(v),f(s))\right)>v_{p_{0}}(n).$ (7)
Hence, using equation 7, we get $v_{p_{0}}(f(u))<\acute{v}_{p_{0}}(n)$ and $v_{p_{0}}(f(v))\geq\acute{v}_{p_{0}}(n)$ for some prime number $p_{0}\mid t$. ∎
4.3 More Necessary and Sufficient Conditions for Equality of Degrees
The following theorem is a generalization of one direction of Theorems 4.3.
Theorem 4.5.
Suppose $D_{n}$ is the maximal Diophantine graph of order $n$ with a labeling $f$ and $u,v\in V(D_{n})$ such that $\deg(u)<n-1$, $\deg(v)<n-1$. If $\omega\big{(}f^{*}(u)\big{)}=\omega\big{(}f^{*}(v)\big{)}$ and for every prime number $p\mid f^{*}(u)$, there exits a prime number $q\mid f^{*}(v)$ such that $\left\lfloor\frac{n}{p^{\acute{v}_{p}(n)}}\right\rfloor=\left\lfloor\frac{n}{q^{\acute{v}_{q}(n)}}\right\rfloor$, then $\deg(u)=\deg(v)$, where $w(n)$ is the number of prime factors of $n$.
Proof.
Suppose $D_{n}$ is the maximal Diophantine graph of order $n$ with a labeling $f$ and $u,v\in V(D_{n})$ such that
$\deg(u)<n-1,\quad\deg(v)<n-1\quad\mbox{and}\quad\omega\big{(}f^{*}(u)\big{)}=\omega\big{(}f^{*}(v)\big{)}.$
Therefore, $\omega\big{(}f^{*}(u)\big{)}=\omega\big{(}f^{*}(v)\big{)}>0.$ Hence, there is a prime number $p$ such that $p\mid f^{*}(u)$. Consequently, our hypothesis implies that there is a prime number $q\mid f^{*}(v)$ such that
$\left\lfloor\frac{n}{p^{\acute{v}_{p}(n)}}\right\rfloor=\left\lfloor\frac{n}{q^{\acute{v}_{q}(n)}}\right\rfloor.$ (8)
Therefore, using Lemmas 2.2 part $(i)$, $p^{\acute{v}_{p}(n)}\mid f(u)$ and $q^{\acute{v}_{q}(n)}\mid f(v).$ Then, using the fundamental theorem of arithmetic, there exist $\grave{u},\grave{v}\in V(D_{n})$ such that
$f(\grave{u}):=\prod\limits_{p\mid f^{*}(u)}p^{\acute{v}_{p}(n)}\ \Big{|}\ f(u)\quad\mbox{and}\quad f(\grave{v}):=\prod\limits_{q\mid f^{*}(v)}q^{\acute{v}_{q}(n)}\ \Big{|}\ f(v).$
By definition of $f(\grave{u})$ and $f(\grave{v})$, we have $f^{*}(u),f^{*}(\grave{u})$ have the same prime factors and $f^{*}(v),f^{*}(\grave{v})$ have the same prime factors. Since $\omega\big{(}f^{*}(u)\big{)}=\omega\big{(}f^{*}(v)\big{)},$ therefore
$\omega\big{(}f^{*}(\grave{u})\big{)}=\omega\big{(}f^{*}(u)\big{)}=\omega\big{(}f^{*}(v)\big{)}=\omega\big{(}f^{*}(\grave{v})\big{)}.$ (9)
Moreover, for every prime number $p$, we have
$p^{\acute{v}_{p}(n)}\mid f(u)\ \mbox{if and only if}\ p^{\acute{v}_{p}(n)}\mid f(\grave{u}).$
Therefore,
$\bigcup\limits_{p|f^{*}(u)}M_{p^{\acute{v}_{p}(n)}}=\bigcup\limits_{p|f^{*}(\grave{u})}M_{p^{\acute{v}_{p}(n)}}.$
Consequently, using Lemma 3.1, for each vertex $s\in V(D_{n})-\{u,\ \grave{u}\}$, we have
$su\in E(D_{n})\quad\mbox{if and only if}\quad s\grave{u}\in E(D_{n}).$
Hence,
$\deg(u)=\deg(\grave{u}).$ (10)
Similarly, we get
$\deg(v)=\deg(\grave{v}).$ (11)
Let $d=(f(\grave{u}),f(\grave{v}))$. Thus, Lemmas 2.2 part $(ii)$ implies that
$\frac{d}{(d,n)}\ \bigg{|}\ f^{*}(\grave{u}),f^{*}(\grave{v}),$ (12)
and hence
$\omega\left(\frac{d}{(d,n)}\right)\leq\omega\big{(}f^{*}(\grave{u})\big{)}=\omega\big{(}f^{*}(\grave{v})\big{)}.$
Let us consider the following cases.
Case 1: Let $\omega\big{(}f^{*}(\grave{u})\big{)}=\omega\big{(}f^{*}(\grave{v})\big{)}=\omega\left(\frac{d}{(d,n)}\right).$ Then, using equation (12), $\frac{d}{(d,n)}$, $f^{*}(\grave{u})$ and $f^{*}(\grave{v})$ have the same prime factors. A similar argument used to prove equation (10) implies that
$\deg(\grave{u})=\deg(\grave{v}).$
Hence, using equations (10) and (11), we get
$\deg(u)=\deg(v).$
Case 2: Let $\omega\big{(}f^{*}(\grave{u})\big{)}=\omega\big{(}f^{*}(\grave{v})\big{)}=\omega\left(\frac{d}{(d,n)}\right)+1=r.$ Then we get $f^{*}(\grave{u})$ and $f^{*}(\grave{v})$ differ in exactly one prime factor. Therefore, using the fundamental theorem of arithmetic, we obtain the following two equations
$f(\grave{u})=\prod\limits_{i=1}^{r-1}h_{i}^{\acute{v}_{h_{i}}(n)}p^{\acute{v}_{p}(n)}\quad\mbox{and}\quad f(\grave{v})=\prod\limits_{i=1}^{r-1}h_{i}^{\acute{v}_{h_{i}}(n)}q^{\acute{v}_{q}(n)},$
where $h_{i},p$ and $q$ are distinct prime numbers such that $h_{i}\mid d,p\nmid d$ and $q\nmid d$ for all $i=1,2,\dots,r-1$. Then, using Theorem 4.2, we have the following two equations
$\begin{split}\deg(\grave{u})=n-\sum_{1\leq i\leq r-1}\left\lfloor\frac{n}{h_{i}^{\acute{v}_{h_{i}}(n)}}\right\rfloor&-\left\lfloor\frac{n}{p^{\acute{v}_{p}(n)}}\right\rfloor+\sum_{1\leq i<j\leq r-1}\left\lfloor\frac{n}{h_{i}^{\acute{v}_{h_{i}}(n)}h_{j}^{\acute{v}_{h_{j}}(n)}}\right\rfloor+\sum_{1\leq i\leq r-1}\left\lfloor\frac{n}{h_{i}^{\acute{v}_{h_{i}}(n)}p^{\acute{v}_{p}(n)}}\right\rfloor-\\ &-\dots+(-1)^{r}\left\lfloor\frac{n}{\prod\limits_{1\leq i\leq r-1}h_{i}^{\acute{v}_{h_{i}}(n)}p^{\acute{v}_{p}(n)}}\right\rfloor,\end{split}$ (13)
and
$\begin{split}\deg(\grave{v})=n-\sum_{1\leq i\leq r-1}\left\lfloor\frac{n}{h_{i}^{\acute{v}_{h_{i}}(n)}}\right\rfloor&-\left\lfloor\frac{n}{q^{\acute{v}_{q}(n)}}\right\rfloor+\sum_{1\leq i<j\leq r-1}\left\lfloor\frac{n}{h_{i}^{\acute{v}_{h_{i}}(n)}h_{j}^{\acute{v}_{h_{j}}(n)}}\right\rfloor+\sum_{1\leq i\leq r-1}\left\lfloor\frac{n}{h_{i}^{\acute{v}_{h_{i}}(n)}q^{\acute{v}_{q}(n)}}\right\rfloor-\\ &-\dots+(-1)^{r}\left\lfloor\frac{n}{\prod\limits_{1\leq i\leq r-1}h_{i}^{\acute{v}_{h_{i}}(n)}q^{\acute{v}_{q}(n)}}\right\rfloor.\end{split}$ (14)
Using Lemma 2.6 and equations (8), (13) and (14), we get
$\deg(\grave{u})=\deg(\grave{v}),$
and hence
$\deg(u)=\deg(v).$
Case 3: Let $\omega\big{(}f^{*}(\grave{u})\big{)}=\omega\big{(}f^{*}(\grave{v})\big{)}=\omega\left(\frac{d}{(d,n)}\right)+2=r.$ Then $f^{*}(\grave{u})$ and $f^{*}(\grave{v})$ differ in exactly two prime factors. Therefore, using the fundamental theorem of arithmetic, we have the following two equations
$f(\grave{u})=\prod\limits_{i=1}^{r-2}h_{i}^{\acute{v}_{h_{i}}(n)}p_{1}^{\acute{v}_{p_{1}}(n)}p_{2}^{\acute{v}_{p_{2}}(n)}\quad\mbox{and}\quad f(\grave{v})=\prod\limits_{i=1}^{r-2}h_{i}^{\acute{v}_{h_{i}}(n)}q_{1}^{\acute{v}_{q_{1}}(n)}q_{2}^{\acute{v}_{q_{1}}(n)},$
where $h_{i},p_{1},p_{2},q_{1}$ and $q_{2}$ are distinct prime numbers satisfying $h_{i}\mid d$, $p_{1}\nmid d$, $p_{2}\nmid d$, $q_{1}\nmid d$ and $q_{2}\nmid d$ for all $i=1,2,\dots,r-2$ such that
$\left\lfloor\frac{n}{p_{1}^{\acute{v}_{p_{1}}(n)}}\right\rfloor=\left\lfloor\frac{n}{q_{1}^{\acute{v}_{q_{1}}(n)}}\right\rfloor\quad\mbox{and}\quad\left\lfloor\frac{n}{p_{2}^{\acute{v}_{p_{2}}(n)}}\right\rfloor=\left\lfloor\frac{n}{q_{2}^{\acute{v}_{q_{2}}(n)}}\right\rfloor.$
Define
$a:=p_{1}^{\acute{v}_{p_{1}}(n)}p_{2}^{\acute{v}_{p_{2}}(n)}\quad\mbox{and}\quad b:=q_{1}^{\acute{v}_{q_{1}}(n)}q_{2}^{\acute{v}_{q_{1}}(n)}.$
Since
$\frac{a}{(a,n)}=p_{1}p_{2}\quad\mbox{and}\quad\frac{b}{(b,n)}=q_{1}q_{2},$
therefore Lemma 2.5 implies that $a>n\ \mbox{or}\ b>n.$ Hence,
$f(\grave{u})\geq a>n\quad\mbox{or}\quad f(\grave{v})\geq b>n,$
which is absurd. The cases $\omega\big{(}f^{*}(\grave{u})\big{)}=\omega\big{(}f^{*}(\grave{v})\big{)}=\omega\left(\frac{d}{(d,n)}\right)+i$, for every $i\geq 3$ can be treated in a very similar way of case $3$ which are absurd. Hence the result follows. ∎
Example 4.3.
In Example 4.2, let $f(u)=14$ and $f(v)=16$, where $u,v\in V(D_{20})$, so $f^{*}(u)=7$ and $f^{*}(v)=8$. Then, applying Theorem 4.5, we have $\omega\left(7\right)=\omega\left(8\right)=1$ and $\left\lfloor\frac{20}{7}\right\rfloor=\left\lfloor\frac{20}{8}\right\rfloor=2.$ Hence, $\deg(u)=\deg(v)=18$, (by applying Lemma 4.1).
The converse of Theorem 4.5 does not hold. For instance, let $u,v\in V(D_{23})$ such that $f(u)=10$, $f(v)=14$. By Theorem 4.2, we have $deg(u)=deg(v)=10$. However, it can be verified that $\left\lfloor\frac{23}{5}\right\rfloor\neq\left\lfloor\frac{23}{7}\right\rfloor$, which demonstrates that the equality of degrees does not imply the equality of values of the floor function.
Corollary 4.3.
Suppose $D_{n}$ is the maximal Diophantine graph of order $n$ with a labeling $f$. Let $u,v\in V(D_{n})$ such that $f^{*}(u)>1$, $f^{*}(v)>1$. If $f^{*}(u),f^{*}(v)$ have the same prime factors, then $\deg(u)=\deg(v)$.
Corollary 4.4.
Suppose $D_{n}$ is the maximal Diophantine graph of order $n$ with a labeling $f$. Let $u,v\in V(D_{n})$ such that $f(v)=tf(u)$ for some $t\geq 1$. If $t\mid n$ and $(t,f(u))=1$, then $\deg(u)=\deg(v)$.
Proof.
Suppose $D_{n}$ is the maximal Diophantine graph of order $n$ with a labeling $f$. Let $u,v\in V(D_{n})$ such that $f(v)=tf(u)$ for some $t\geq 1$, $t\mid n$ and $(t,f(u))=1$. The reduced labels $f^{*}(v),f^{*}(u)$ satisfy the following equalities.
$f^{*}(v)=\frac{f(v)}{(f(v),n)}=\frac{tf(u)}{(tf(u),n)}=\frac{tf(u)}{t(f(u),n)}=\frac{f(u)}{(f(u),n)}=f^{*}(u).$
Therefore, the two reduced labels $f^{*}(u),f^{*}(v)$ have the same primes factors or each of them is equal to 1. Hence, using Corollary 4.3 and Theorem 4.1, the result follows. ∎
The following two examples shows the converse of Corollary 4.3 is not true,
Example 4.4.
Suppose $D_{n}$ is the maximal Diophantine graph of order $n$ with a labeling $f$. Let $u,v\in V(D_{n})$ such that $f(u)\neq f(v)$, $f(u)\mid f(v)$, and $f(v)=p^{\acute{v}_{p}(n)}$, $\frac{n}{2}<f(v)<n$, where $p$ is a prime number, using Theorem 4.1, we have
$\deg(u)=n-1=\deg(v).$
However, $f^{*}(u)=\frac{f(u)}{(f(u),n)}=1$ and $f^{*}(v)=\frac{p^{\acute{v}_{p}(n)}}{\left(p^{\acute{v}_{p}(n)},n\right)}=p$ have not the same prime factors.
Example 4.5.
Suppose $D_{n}$ is the maximal Diophantine graph of even order $n$ with a labeling $f$. Let $u,v\in V(D_{n})$ such that $f(u)=2p^{\acute{v}_{p}(n)}$ and $f(v)=2q^{\acute{v}_{q}(n)}$, $\frac{n}{3}<p^{\acute{v}_{p}(n)},q^{\acute{v}_{q}(n)}<\frac{n}{2}$, where $p,q$ are distinguish prime numbers, this means $f(u)\nmid f(v)$, using Lemma 4.1 and Lemmas 2.2 part $(i)$, we have
$\deg(u)=n-\left\lfloor\frac{n}{p^{\acute{v}_{p}(n)}}\right\rfloor=n-2=n-\left\lfloor\frac{n}{q^{\acute{v}_{q}(n)}}\right\rfloor=\deg(v).$
However, $f^{*}(u)=p$ and $f^{*}(v)=q$ have distinct prime factors.
The last two examples motivate the following partial converse of Corollary 4.3.
Theorem 4.6.
Suppose $D_{n}$ is the maximal Diophantine graph of order $n$ with a labeling $f$. Let $u,v\in V(D_{n})$ such that $f(u)\mid f(v)$, $f(v)$ is not a prime power number and $f^{*}(u)>1$. If $\deg(u)=\deg(v)$, then $f^{*}(u),f^{*}(v)$ have the same prime factors.
Proof.
Suppose $D_{n}$ is the maximal Diophantine graph of order $n$ with a labeling $f$. Let $u,v\in V(D_{n})$ such that $f(u)\mid f(v)$, $f(v)$ is not a prime power number, $\deg(u)=\deg(v)$ and $f^{*}(u)>1$. Then, using Lemmas 2.2 part $(ii)$, we have $f^{*}(u)\mid f^{*}(v)$. Consequently, $f^{*}(v)>1$ and also any prime factor of $f^{*}(u)$ is a prime factor of $f^{*}(v)$. Let there exists a prime number $p$ such that $p\mid f^{*}(v)$. Then, using Lemmas 2.2 part $(i)$, $p^{\acute{v}_{p}(n)}\mid f(v)$, i.e., $f(v)\in M_{p^{\acute{v}_{p}(n)}}$. Let a vertex $s\in V(D_{n})$ such that $f(s)=p^{\acute{v}_{p}(n)}\in M_{p^{\acute{v}_{p}(n)}}.$ Since $f(v)$ is not a prime power number, we get $s\neq v$. Using Lemma 3.1, we have $(f(v),f(s))\nmid n$. Therefore, $s\notin N(v)$. Consequently, using Lemma 4.3, we obtain $s\notin N(u)$. Then we get $\left(f(u),p^{\acute{v}_{p}(n)}\right)\nmid n$. Thus, $p^{\acute{v}_{p}(n)}\mid f(u)$. Consequently, using Lemmas 2.2 part $(i)$, $p\mid f^{*}(u)$. Therefore, for every prime number $p$,
$p\mid f^{*}(u)\quad\mbox{if and only if}\quad p\mid f^{*}(v).$
Hence, the two reduced labels $f^{*}(u),f^{*}(v)$ have the same prime factors. ∎
Future Work
Give two efficient algorithms that one determines the maximal Diophantine graphs and anther determines whether a given graph is Diophantine or not. Furthermore, exploring different variants of Diophantine labeling can further enrich the theory and uncover new connections between graph theory and number theory.
Declaration
The authors contributed equally to the manuscripts and approved the submitted scripts. We declare that we have no conflicts of interests. We confirm that this work is original and has not been published elsewhere, nor is it currently under consideration for publication elsewhere.

### v2

@@
-Diophantine Graphs
+Linear Diophantine Graphs
 Abstract
-This manuscript introduces Diophantine labeling, a new way of labeling of the vertices for finite simple undirected graphs with some divisibility condition on the edges.
-Maximal graphs admitting Diophantine labeling are investigated and their number of edges are computed.
-Some number-theoretic techniques are used to characterize vertices of maximum degree and nonadjacent vertices.
-Some necessary and sufficient conditions for vertices of equal degrees are found.
-Keywords: Graph labeling, Prime graph, Diophantine graph, $p$-adic valuation.
-MSC code: 05A10, 05C07, 05C78, 11A05, 11A25, 11B75, 11D04, 11D88.
+This manuscript introduces linear Diophantine labelling, a new method for assigning labels to the vertices of finite, simple, undirected graphs.
+A key feature of this method is a divisibility condition imposed on the edges, incorporating number-theoretic properties into graph labelling.
+The study focuses on identifying maximal graphs that admit such labellings and computes their number of edges and the degree of each vertex.
+Number-theoretic techniques are employed to examine structural properties, including the characterization of maximum degree vertices and conditions for nonadjacent vertices.
+The manuscript also establishes necessary and sufficient conditions for vertices with equal degrees, offering new insights into the interaction between graph theory and number theory.
+Keywords: Graph labelling, Prime graph, Diophantine graph, Equal degrees, $p$-adic valuation.
+AMS Subject Classification: 05C07, 05C78, 11A05, 11A25, 11D88.
 1 Introduction
-A graph labeling is an assignment of real values to the vertices or edges that meet some conditions.
-The definition of prime labeling was originated by R.
-Entringer and introduced formally in a paper by R.
+A graph labelling is an assignment of real values to the vertices or edges or both edges and vertices that meet some conditions.
+It constitutes a significant area within graph theory and has numerous applications, including in coding theory, radar systems, astronomy, missile guidance, communication networks, cryptography, network security, X-ray crystallography, and database management.
+In this manuscript, we investigate a new type of labeling, referred to as linear Diophantine labeling, which constitutes a generalization of prime labeling.
+Labeling problems are pivotal in studying computational complexity and algorithmic graph theory, serving as a bridge between number theory and graph theory.
+The concept of prime labelling was first introduced by R.
+Entringer and later formally defined in a paper by R.
 Tout [12].
-A graph $G=(V,E)$ with $n$ vertices has a prime labeling if there exists a bijective map $f:V\rightarrow\{1,2,\dots,n\}$ such that for every $uv\in E$, $(f(u),f(v))=1$.
-Many results of prime labeling and other types of prime labeling can be found in the dynamic survey of graph labeling by J.A.
+A simple graph $G=(V,E)$ with $n$ vertices is said to have a prime labelling if there is a bijective function $f$ mapping $V$ to $\{1,2,\dots,n\}$ in which for every $uv\in E$, the values $f(u),f(v)$ are relatively prime, where $V$ is the vertex set and $E$ is the edge set.
+Many results of prime labelling and other types of prime labelling can be found in the dynamic survey of graph labelling by J.A.
 Gallian [6].
-Other variants of prime labeling have been introduced by many researchers (see for example [4], [5], [8], [13]).
+Other variants of prime labelling have been introduced by many researchers (see, for example, [4], [5], [8], [14]).
 This paper gives a generalization for the concept of prime graphs.
-More precisely, we define a graph $G$ with $n$ vertices to be a Diophantine graph if there exists a bijective map $f:V\rightarrow\{1,2,\dots,n\}$ such that for every $uv\in E$, $(f(u),f(v))\mid n$.
+More precisely, a simple graph $G$ with $n$ vertices is defined as a linear Diophantine graph (or simply, a Diophantine graph) if there is a bijective function $f$ mapping $V$ to $\{1,2,\dots,n\}$ in which for every $uv\in E$, the greatest common divisor of $f(u)$ and $f(v)$ divides $n$.
 Clearly, any prime graph is a Diophantine graph.
 Moreover, any Diophantine graph with a prime number of vertices is a prime graph.
-In the maximal Diophantine graph with $n$ vertices, an explicit formulas of the number of edges and the degree of a vertex are proved, the full degree vertices are determined.
-Furthermore, we provide sufficient and necessary conditions that determine the equality of degrees among non-full degree vertices in the maximal Diophantine graphs.
+In a maximal Diophantine graph with $n$ vertices, an explicit formula of the number of edges and the degree of a vertex are proved, the full degree vertices are determined.
+Furthermore, we provide sufficient and necessary conditions that determine the equality of degrees among vertices whose degrees are smaller than n1n-1 in a maximal Diophantine graphs.
 The paper is organized as follows.
 Section 2 includes some definitions, preliminaries, lemmas and direct observations.
-Section 3 presents the number of edges formula in the maximal Diophantine graph with $n$ vertices.
-Section 4 is partitioned into three subsections, each unveiling important results related to the maximal Diophantine graph with order $n$.
-Subsection 4.1 covers a characterization of labels of vertices with full degree and a formula that compute the degrees of any vertex.
+Section 3 presents the number of edges formula in a maximal Diophantine graph with $n$ vertices.
+Section 4 is split into three subsections, each unveiling important results related to a maximal Diophantine graph with order $n$.
+Subsection 4.1 covers a characterization of labels of vertices with full degree and a formula that computes the degrees of any vertex.
 Subsection 4.2 and 4.3 encompass some necessary and sufficient conditions for equality of degrees.
 For definitions and terminologies of graph theory, we follow F. Harary [7] and A. Bickle [2]. Also, we follow D. Burton [3] and K.H. Rosen [9] for basic definitions and notations in number theory.
+Theorem 1.1.
+[3], [9] (Fundamental Theorem of Arithmetic)
+Every positive integer $n>1$ can be expressed as a product of primes; this representation is unique, apart from the order in which the factors occur.
+Definition 1.1.
+[2] Two graphs $G_{1},G_{2}$ are said to be isomorphic, denoted by $G_{1}\cong G_{2}$, if there exists a bijective map $\varphi:V(G_{1})\rightarrow V(G_{2})$ such that for all $u,v\in V(G_{1})$,
+$uv\in E(G_{1})\quad\mbox{if and only if}\quad\varphi(u)\varphi(v)\in E(G_{2}).$
 2 Definitions and Preliminaries
-In the beginning of this section, we give the definitions of the prime graphs and Diophantine graphs. It also explores the relation between them and some direct observations.
+In the beginning of this section, we present the definitions of the prime graphs and the Diophantine graphs and we explore the relation between them.
+Additionally, we provide direct observations on several number-theoretic lemmas.
 Definition 2.1.
-[11], [12] Let $G$ be a graph with $n$ vertices.
-The graph $G$ is called a prime graph if there exists a bijective map $f:V\rightarrow\{1,2,\dots,n\}$ such that for all $uv\in E$, $(f(u),f(v))=1$.
-Such a map $f$ is called a prime labeling of $G$.
-The maximal prime graph with $n$ vertices, denoted by $R_{n}$, is a prime graph such that adding any new edge yields a non-prime graph.
+[11], [12] A simple graph $G$ of order $n$ is called a prime graph if there exists a bijective mapping $f:V\rightarrow\{1,2,\dots,n\}$ such that $(f(u),f(v))=1$ for all $uv\in E$.
+The function $f$ is called a prime labelling of $G$.
+A prime graph of order $n$ is called a maximal prime graph, denoted by $(R_{n},f)$, if any new edge added to this graph turns it into a non-prime graph.
 Definition 2.2.
-Let $G$ be a graph with $n$ vertices.
-The graph $G$ is called a Diophantine graph if there exists a bijective map $f:V\rightarrow\{1,2,\dots,n\}$ such that for all $uv\in E$, $(f(u),f(v))\mid n$.
-Such a map $f$ is called a Diophantine labeling of $G$.
-The maximal Diophantine graph with $n$ vertices, denoted by $D_{n}$, is a Diophantine graph such that adding any new edge yields a non-Diophantine graph.
+A simple graph $G$ of order $n$ is called a linear Diophantine graph (or simply, a Diophantine graph) if there exists a bijective mapping $f:V\rightarrow\{1,2,\dots,n\}$ such that $(f(u),f(v))\mid n$ for all $uv\in E$.
+The function $f$ is called a Diophantine labelling of $G$.
+A Diophantine graph of order $n$ is called a maximal Diophantine graph, denoted by $(D_{n},f)$, if any new edge added to this graph turns it into a non-Diophantine graph.
+If there is no ambiguity, we omit $f$ from $(D_{n},f)$ and simply write $D_{n}$.
 Example 2.1.
-The Petersen graph $P$ is a Diophantine graph, but the complete graph $K_{5}$ is not.
-66662222333344441010101011117777888899995555
-11112222333344445555
+The Peterson graph $P$ is a Diophantine graph, but the complete graph $K_{5}$ is not.
+6622334410101177889955
+1122334455
 Figure 1: $P$ is a Diophantine graph while $K_{5}$ is not
@@
 The following are some maximal Diophantine graphs.
-1111222233334444
-11112222333344445555
-111122223333444455556666
-1111555577773333222266664444
-11112222333344445555666677778888
+11223344
+1122334455
+112233445566
+11557733226644
+1122334455667788
 Figure 2: Maximal Diophantine Graphs $D_{4},\dots,D_{8}$
 It is easy to see that every prime graph is a Diophantine graph, but the converse is not necessarily true, e.g., $D_{4},D_{6},D_{8}$.
-However, in the case that $n$ is a prime number, a graph is Diophantine if and only if it is a prime graph.
-Moreover, $D_{n}\cong R_{n}.$ The proof is straightforward using the following lemma.
+Theorem 2.1.
+Let $n>1$ be a prime number. A graph $G$ of order $n$ is a Diophantine graph if and only if $G$ is a prime graph.
+Proof.
+The proof is clear. ∎
+Theorem 2.2.
+For every $n>1$, $D_{n}\cong R_{n}$ if and only if $n$ is a prime number.
+Proof.
+Let $K_{n}$ be the complete graph of order $n$ and $f:V(K_{n})\rightarrow\{1,2,\dots,n\}$ be a bijective map. Define two sets
+$E(R_{n}):=\{uv\in E(K_{n}):(f(u),f(v))=1\}$
+and
+$E(D_{n}):=\{uv\in E(K_{n}):(f(u),f(v))\mid n\}.$
+Let $D_{n}\cong R_{n}$. Then $E(R_{n})=E(D_{n})$. Therefore,
+$(f(u),f(v))\mid n\quad\mbox{if and only if}\quad(f(u),f(v))=1.$
+Suppose, by the way of contradiction, $n>1$ is a composite number.
+Then there exists a prime number $p\mid n$ such that $p\leq\frac{n}{2}$.
+Let $f(u)=p\leq n$ and $f(v)=2p\leq n$.
+Then $(f(u),f(v))=p\mid n$ and $(f(u),f(v))\neq 1$ which contradicts $(f(u),f(v))=1$.
+Hence, $n$ is a prime number.
+Conversely, let $n$ be a prime number and let $uv\in E(R_{n})$.
+Then $(f(u),f(v))=1$.
+Consequently, $(f(u),f(v))\mid n$.
+Therefore, $uv\in E(D_{n})$.
+On the other hand, let $uv\in E(D_{n})$.
+Then $(f(u),f(v))\mid n$.
+Since $n$ is a prime number, we have either $(f(u),f(v))=1$ or $(f(u),f(v))=n$.
+In the case that $(f(u),f(v))=1$, we have nothing to proof.
+In the other case that $(f(u),f(v))=n,$ we get $f(u)=f(v)=n$ which is a contradiction with the assumption.
+Consequently, $(f(u),f(v))=1$.
+Therefore, $uv\in E(R_{n})$.
+Hence, $D_{n}\cong R_{n}$.
+∎
+In the subsequent sections, the $p$-adic valuation of $n\in\mathbb{Z}^{+}$ is defined by $v_{p}(n):=t$ such that $p^{t}\mid n$ and $p^{t+1}\nmid n$, where $p$ is a prime number and $t\in\mathbb{N}:=\mathbb{Z}^{+}\cup\{0\}$ and $\mathbb{Z}^{+}$ is the set of positive integers.
+The set of all multiple of $a\in\mathbb{Z}^{+}$ up to $n$ is denoted by $M_{a}$, i.e.
+$M_{a}:=\left\{ka\ :\ k=1,2,\dots,\left\lfloor\frac{n}{a}\right\rfloor\right\}.$ In any graph $G$ of order $n$, we call $f^{*}(u):=\frac{f(u)}{(f(u),n)}$ the reduced label of a vertex $u$ and the neighborhood of $s\in V(G)$ is defined by $N(s):=\{u\in V(G):su\in E(G)\}$.
+Now, we present a crucial definition and important number-theoretic lemmas that will allow us to formulate and prove our results smoothly.
+Definition 2.3.
+A number $p^{\acute{v}_{p}(n)}$ is called a critical prime power number with respect to a prime number $p$ and a positive integer $n$, where $\acute{v}_{p}(n):=v_{p}(n)+1$ is called the successor of the $p$-adic valuation.
 Lemma 2.1.
-Let $n>1$. Then we have $n$ is a prime number  if and only if  for all $a,b\leq n$, $a\neq b$, [$(a,b)=1$ if and only if $(a,b)\mid n$].
-Proof.
-Let $n$ be a prime number, $a,b<n$ and $a\neq b$.
-If $(a,b)=1,$ then $(a,b)\mid n$.
-On the other hand, if $(a,b)\mid n,$ then either $(a,b)=1$ or $(a,b)=n$.
-In the case that $(a,b)=1$, we have nothing to proof.
-In the other case that $(a,b)=n,$ we get $a=b=n$ which is a contradiction with the hypothesis.
-Conversely, Assume $a\neq b$ and [$(a,b)=1$ if and only if $(a,b)\mid n$] for all $a,b<n$.
-Suppose, by the way of contradiction, $n>1$ is a composite number.
-Then there exists prime $p\mid n$ such that $p\leq\frac{n}{2}$.
-Let $a=p\leq n$, $b=2p\leq n$.
-Hence, $(a,b)=p\mid n$ and $(a,b)\neq 1,$ which contradicts the assumption.
-Hence, the result follows.
-∎
-In the subsequent sections, in any graph $G$ of order $n$, we call $f^{*}(u):=\frac{f(u)}{(f(u),n)}$ the reduced label of a vertex $u$ and we denote the set of all multiples of a positive integer $a$ up to $n$ by $M_{a}$, i.e.
-$M_{a}:=\{ta\ :\ t=1,2,\dots,\lfloor\frac{n}{a}\rfloor\}.$ Now, we present a crucial definition and an important number-theoretic lemmas that will allow us to formulate and prove our results smoothly.
-Definition 2.3.
-For a given positive integer $n$ and prime number $p$, the successor of the $p$-adic valuation is defined by $\acute{v}_{p}(n):=v_{p}(n)+1,$ where $v_{p}(n)$ is the $p$-adic valuation of an integer $n.$ The number $p^{\acute{v}_{p}(n)}$ is called the critical prime power number with respect to $p,n$.
-Lemma 2.2.
-Let $u,v\in V(D_{n})$ and $f$ be a Diophantine labeling of $D_{n}$. Then we have
-  • (i)
+Let $u,v$ be two vertices in $D_{n}$ and $f$ be a Diophantine labelling of $D_{n}$. Then
+  1. i.
     for every prime number $p$, $p\mid f^{*}(u)$ if and only if $p^{\acute{v}_{p}(n)}\mid f(u);$
-  • (ii)
+  2. ii.
     if $f(u)\mid f(v)$, then $f^{*}(u)\mid f^{*}(v).$
 Proof.
-The proof of part $(i)$ is straightforward. Let us prove the second part. The $p$-adic valuation function [1] is extended normally to rational numbers as follows
+The proof of part $i.$ is straightforward. Let us prove the second part. The $p$-adic valuation function [1] is extended normally to rational numbers as follows
 $v_{p}\left(\frac{b}{a}\right):=v_{p}(b)-v_{p}(a),$
-for any $a,b\in\mathbb{Z}^{+}$ and for every prime number $p.$ Let $u,v\in V(D_{n})$ such that $f(u)\mid f(v).$ Using the basic properties of the $p$-adic valuation function, we have
+for any $a,b\in\mathbb{Z}^{+}$ and for every prime number $p$. Let $u,v$ be two vertices in $V(D_{n})$ such that $f(u)\mid f(v).$ Using the basic properties of the $p$-adic valuation function, we have
 $\begin{split}v_{p}\left(\frac{f(v)}{f(u)}\frac{(f(u),n)}{(f(v),n)}\right)&=v_{p}(f(v))-v_{p}(f(u))+v_{p}\big((f(u),n)\big)-v_{p}\big((f(v),n)\big).\\ &=v_{p}(f(v))-v_{p}(f(u))+\min\big\{v_{p}(f(u)),v_{p}(n)\big\}-\min\big\{v_{p}(f(v)),v_{p}(n)\big\}.\end{split}$
@@
 in all possible cases between $v_{p}(f(u)),v_{p}(n)$ and $v_{p}(f(v))$, namely
-$v_{p}(n)<v_{p}(f(u)),\quad v_{p}(f(u))\leq v_{p}(n)\leq v_{p}(f(v))\quad\mbox{and}\quad v_{p}(f(v))<v_{p}(n).$
-Hence,
-$\frac{f(u)}{(f(u),n)}\ \Bigg{|}\ \frac{f(v)}{(f(v),n)},$
-which completes the proof. ∎
-In the proof of Lemma 2.3, we will use the following identity: $\frac{1}{j}-\frac{1}{j+1}=\left(1+\frac{2}{j}\right)\left[\frac{1}{j+1}-\frac{1}{j+2}\right]$, $j\in\mathbb{Z}^{+}$.
+$v_{p}(n)<v_{p}(f(u)),\ v_{p}(f(u))\leq v_{p}(n)\leq v_{p}(f(v))\ \mbox{and}\ v_{p}(f(v))<v_{p}(n).$
+Therefore,
+$\frac{f(u)}{(f(u),n)}\ \Bigg|\ \frac{f(v)}{(f(v),n)}.$
+Hence, the proof follows. ∎
+We will use the following identity: $\frac{1}{j}-\frac{1}{j+1}=\left(1+\frac{2}{j}\right)\left[\frac{1}{j+1}-\frac{1}{j+2}\right]$, $j\in\mathbb{Z}^{+}$ for the proof of Lemma 2.2.
 Lemma 2.2.
@@
 Clearly, we have
-$\begin{cases}\lfloor\sqrt{n}\rfloor\in\left(\frac{n}{\lfloor\sqrt{n}\rfloor+1},\frac{n}{\lfloor\sqrt{n}\rfloor}\right]\quad\mbox{and}\quad\frac{n}{\lfloor\sqrt{n}\rfloor}-\frac{n}{\lfloor\sqrt{n}\rfloor+1}<1,&\quad\mbox{if}\quad n<\lfloor\sqrt{n}\rfloor(\lfloor\sqrt{n}\rfloor+1);\\ &\\ \lfloor\sqrt{n}\rfloor\notin\left(\frac{n}{\lfloor\sqrt{n}\rfloor+1},\frac{n}{\lfloor\sqrt{n}\rfloor}\right]\quad\mbox{and}\quad\frac{n}{\lfloor\sqrt{n}\rfloor}-\frac{n}{\lfloor\sqrt{n}\rfloor+1}\geq 1,&\quad\mbox{if}\quad n\geq\lfloor\sqrt{n}\rfloor(\lfloor\sqrt{n}\rfloor+1).\\ \end{cases}$
+[…]
 Since $\lfloor\sqrt{n}\rfloor<\frac{n}{\lfloor\sqrt{n}\rfloor}$ and $\frac{n}{j}-\frac{n}{j+1}>\frac{n}{j+1}-\frac{n}{j+2},$ therefore there exists $j\geq\lfloor\sqrt{n}\rfloor$ such that
@@
 Proof.
-Let $a,b$ and $n\in\mathbb{N}$ such that $\lfloor\frac{n}{a}\rfloor=\lfloor\frac{n}{b}\rfloor$.
+Clearly, the statement is true in case of positive integers $a$ or $b$ are greater than or equal to $n\in\mathbb{Z}^{+}$.
+Now, let $a,b<n$ such that $\lfloor\frac{n}{a}\rfloor=\lfloor\frac{n}{b}\rfloor$.
 Therefore, there exists $j\in\{1,2,\dots,n\}$ such that $\frac{n}{j+1}<a,b\leq\frac{n}{j}$.
@@
 $\left\lfloor\frac{n}{p_{1}^{\acute{v}_{p_{1}}(n)}}\right\rfloor=\left\lfloor\frac{n}{q_{1}^{\acute{v}_{q_{1}}(n)}}\right\rfloor\quad\mbox{and}\quad\left\lfloor\frac{n}{p_{2}^{\acute{v}_{p_{2}}(n)}}\right\rfloor=\left\lfloor\frac{n}{q_{2}^{\acute{v}_{q_{2}}(n)}}\right\rfloor.$ (2)
-Since $\frac{a}{(a,n)}=p_{1}^{\alpha_{1}}p_{2}^{\alpha_{2}}$ and $\frac{b}{(b,n)}=q_{1}^{\beta_{1}}q_{2}^{\beta_{2}},$ therefore Lemmas 2.2 part $(i)$ implies that
+Since $\frac{a}{(a,n)}=p_{1}^{\alpha_{1}}p_{2}^{\alpha_{2}}$ and $\frac{b}{(b,n)}=q_{1}^{\beta_{1}}q_{2}^{\beta_{2}},$ therefore Lemmas 2.1, part $i.$, implies that
 $p_{1}^{\acute{v}_{p_{1}}(n)}p_{2}^{\acute{v}_{p_{2}}(n)}\ \Big|\ a\quad\mbox{and}\quad q_{1}^{\acute{v}_{q_{1}}(n)}q_{2}^{\acute{v}_{q_{2}}(n)}\ \Big|\ b.$ (3)
@@
 $n<p_{2}^{\acute{v}_{p_{2}}(n)}q_{2}^{\acute{v}_{q_{2}}(n)}<q_{1}^{\acute{v}_{q_{1}}(n)}q_{2}^{\acute{v}_{q_{2}}(n)}\leq b.$
-Hence, the result follows. ∎
+Hence, the proof follows. ∎
 Lemma 2.5.
 Let $a,b,t,n\in\mathbb{Z}^{+}$. If $\left\lfloor\frac{n}{a}\right\rfloor=\left\lfloor\frac{n}{b}\right\rfloor$, then $\left\lfloor\frac{n}{ta}\right\rfloor=\left\lfloor\frac{n}{tb}\right\rfloor$.
-The proof of Lemma 2.6 is a direct application of the fact $\left\lfloor\frac{x}{n}\right\rfloor=\left\lfloor\frac{\lfloor x\rfloor}{n}\right\rfloor$ for any $x\in\mathbb{R}$ and $n\in\mathbb{Z}^{+}.$ Lemmas 2.5 and 2.6 will be used in the proof of Theorem 4.5.
+The proof of Lemma 2.5 is a direct application of the fact $\left\lfloor\frac{x}{n}\right\rfloor=\left\lfloor\frac{\lfloor x\rfloor}{n}\right\rfloor$ for any $x\in\mathbb{R}$ and $n\in\mathbb{Z}^{+},$ where $\mathbb{R}$ is the set of real numbers.
+Lemmas 2.4 and 2.5 will be used in the proof of Theorem 4.6.
 3 Number of Edges of $D_{n}$
-The following lemma is a characterization of nonadjacent vertices in the maximal Diophantine graphs $D_{n}$.
+In this section, we derive a formula for the number of edges of maximal Diophantine graphs $D_{n}$ using a lemma that characterizes nonadjacent vertices.
+In the remainder of this paper, let $D_{n}$ denote a maximal Diophantine graph with order $n$ that admits a Diophantine labelling $f$.
 Lemma 3.1.
-Suppose $D_{n}$ is the maximal Diophantine graph of order $n$ with a labeling $f$. For each $u,v\in V(D_{n}),$ $uv\notin E(D_{n})$ if and only if there exists a prime number $p$ such that
-$f(u),f(v)\in M_{p^{\acute{v}_{p}(n)}}:=\left\{kp^{\acute{v}_{p}(n)}:\ k=1,2,\dots,\left\lfloor\frac{n}{p^{\acute{v}_{p}(n)}}\right\rfloor\right\}.$
-Proof.
-Suppose $D_{n}$ is the maximal Diophantine graph of order $n$ with a labeling $f$ and let $u,v\in V(D_{n})$. Then we have
-$\begin{split}uv\notin E(D_{n})&\quad\mbox{if and only if}\quad(f(u),f(v))\nmid n\\ &\quad\mbox{if and only if}\quad v_{p}\left(\big{(}f(u),f(v)\big{)}\right)>v_{p}(n)\quad\mbox{for some prime number $p$}\\ &\quad\mbox{if and only if}\quad\min\left\{v_{p}(f(u)),\ v_{p}(f(v))\right\}\geq v_{p}(n)+1=\acute{v}_{p}(n)\quad\mbox{for some prime number $p$}\\ &\quad\mbox{if and only if}\quad v_{p}(f(u))\geq\acute{v}_{p}(n)\quad\mbox{and}\quad v_{p}(f(v))\geq\acute{v}_{p}(n)\quad\mbox{for some prime number $p$}\\ &\quad\mbox{if and only if}\quad p^{\acute{v}_{p}(n)}\mid f(u)\quad\mbox{and}\quad p^{\acute{v}_{p}(n)}\mid f(v)\quad\mbox{for some prime number $p$}\\ &\quad\mbox{if and only if}\quad f(u),\ f(v)\in M_{p^{\acute{v}_{p}(n)}}\quad\mbox{for some prime number $p$.}\hskip 139.41832pt\square\end{split}$
+For each two vertices $u,v$ in $D_{n}$, $uv\notin E(D_{n})$ if and only if there exists a prime number $p$ such that $f(u),f(v)\in M_{p^{\acute{v}_{p}(n)}},$ where $M_{p^{\acute{v}_{p}(n)}}$ is the set of multiple of $p^{\acute{v}_{p}(n)}$ up to $n$.
+Proof. Let $u,v$ be two vertices in $D_{n}$ such that
+[…]
 It is clear that from the Lemma 3.1, if such a prime number $p$ exists, then $p<\frac{n}{2}$.
@@
 Theorem 3.1.
-The number of edges of the maximal Diophantine graph $D_{n}$ with order $n$ is given by
-$C(n,2)-\sum_{1\leq i\leq\pi(n)}C\Bigg{(}\Bigg{\lfloor}\frac{n}{p_{i}^{\acute{v}_{p_{i}}(n)}}\Bigg{\rfloor},2\Bigg{)}+\sum_{1\leq i<j\leq\pi(n)}C\Bigg{(}\Bigg{\lfloor}\frac{n}{p_{i}^{\acute{v}_{p_{i}}}(n)p_{j}^{\acute{v}_{p_{j}}(n)}}\Bigg{\rfloor},2\Bigg{)}-\dots+(-1)^{\pi(n)}C\left(\left\lfloor\frac{n}{\prod\limits_{1\leq i\leq\pi(n)}p_{i}^{\acute{v}_{p_{i}}(n)}}\right\rfloor,2\right),$
-where $p_{1},p_{2},\dots,p_{\pi(n)}<n$ are distinct prime numbers, $\pi(n)$ is the number of prime numbers not exceeding $n$ and $C(n,r)$ is the binomial coefficient with $C(n,r):=0$ if $n<r$ and $C(n,r):=\binom{n}{r}$ if $n\geq r$.
+The size of $D_{n}$ is given by
+$\begin{split}|E(D_{n})|=C(n,2)&-\sum_{1\leq i\leq\pi(n)}C\Bigg(\Bigg\lfloor\frac{n}{p_{i}^{\acute{v}_{p_{i}}(n)}}\Bigg\rfloor,2\Bigg)+\sum_{1\leq i<j\leq\pi(n)}C\Bigg(\Bigg\lfloor\frac{n}{p_{i}^{\acute{v}_{p_{i}}}(n)p_{j}^{\acute{v}_{p_{j}}(n)}}\Bigg\rfloor,2\Bigg)\\ &-\dots+(-1)^{\pi(n)}C\left(\left\lfloor\frac{n}{\prod\limits_{1\leq i\leq\pi(n)}p_{i}^{\acute{v}_{p_{i}}(n)}}\right\rfloor,2\right),\end{split}$
+where $p_{i}<n$, $i=1,2,\dots\pi(n)$ are distinct prime numbers, $\pi(n)$ is the number of prime numbers not exceeding $n$ and $C(n,r)$ is the binomial coefficient, defined as $C(n,r):=0$ if $n<r$ and $C(n,r):=\binom{n}{r}$ if $n\geq r$.
 Proof.
-Let $D_{n}$ be the maximal Diophantine graph with $n$ vertices, $f$ be a Diophantine labeling of $D_{n}$ and $p_{i}<n,\ i=1,2,\dots,\pi(n)$ be distinct prime numbers.
-Using Lemma 3.1, we get for each $u,v\in V(D_{n})$, $uv\notin E(D_{n})$ if and only if $f(u),f(v)\in M_{p_{i}^{\acute{v}_{p_{i}}(n)}}$ for some $i=1,2,\dots,\pi(n)$.
+Let $p_{i}<n$, $i=1,2,\dots,\pi(n)$ be distinct prime numbers.
+Then, using Lemma 3.1, we get for each two vertices $u,v$ in $D_{n}$, $uv\notin E(D_{n})$ if and only if $f(u),f(v)\in M_{p_{i}^{\acute{v}_{p_{i}}(n)}}$ for some $i=1,2,\dots,\pi(n)$.
 Define
-$S_{p_{i}}:=\big{\{}\{u,v\}\subset V(D_{n}):f(u),f(v)\in M_{p_{i}^{\acute{v}_{p_{i}}(n)}},\ f(u)\neq f(v)\big{\}},$
-$i=1,2,\dots,\pi(n)$. So, we have the following equation
+$S_{p_{i}}:=\left\{uv\notin E(D_{n}):f(u),f(v)\in M_{p_{i}^{\acute{v}_{p_{i}}(n)}}\right\},\quad i=1,2,\dots,\pi(n).$
+Thus, we have the following equation
 $\begin{split}|E(D_{n})|&=|E(K_{n})|-|E(\overline{D_{n}})|,\ \mbox{where $\overline{D_{n}}$ is the complement of $D_{n}$.}\\ &=|E(K_{n})|-\left|\bigcup\limits_{1\leq i\leq\pi(n)}S_{p_{i}}\right|.\end{split}$
 Using the inclusion-exclusion principle (see [10]), we have
-$|E(D_{n})|=C(n,2)-\left[\sum_{1\leq i\leq\pi(n)}|S_{p_{i}}|-\sum_{1\leq i<j\leq\pi(n)}|S_{p_{i}}\cap S_{p_{j}}|+\dots+(-1)^{\pi(n)-1}\left|\bigcap\limits_{1\leq i\leq\pi(n)}S_{p_{i}}\right|\right].$
-Therefore,
-$\begin{split}|E(D_{n})|&=C(n,2)-\sum_{1\leq i\leq\pi(n)}C\left(\left|M_{p_{i}^{\acute{v}_{p_{i}}(n)}}\right|,2\right)+\sum_{1\leq i<j\leq\pi(n)}C\left(\left|M_{p_{i}^{\acute{v}_{p_{i}}(n)}}\cap M_{p_{j}^{\acute{v}_{p_{j}}(n)}}\right|,2\right)-\\ &\qquad\qquad\quad\dots+(-1)^{\pi(n)}C\left(\left|\bigcap\limits_{1\leq i\leq\pi(n)}M_{p_{i}^{\acute{v}_{p_{i}}(n)}}\right|,2\right).\\ &=C(n,2)-\sum_{1\leq i\leq\pi(n)}C\left(\left\lfloor\frac{n}{p_{i}^{\acute{v}_{p_{i}}(n)}}\right\rfloor,2\right)+\sum_{1\leq i<j\leq\pi(n)}C\left(\left\lfloor\frac{n}{p_{i}^{\acute{v}_{p_{i}}(n)}p_{j}^{\acute{v}_{p_{j}}(n)}}\right\rfloor,2\right)-\\ &\qquad\qquad\quad\dots+(-1)^{\pi(n)}C\left(\left\lfloor\frac{n}{\prod\limits_{1\leq i\leq\pi(n)}p_{i}^{\acute{v}_{p_{i}}(n)}}\right\rfloor,2\right).\end{split}$
+$\begin{split}|E(D_{n})|&=C(n,2)-\sum_{1\leq i\leq\pi(n)}|S_{p_{i}}|+\sum_{1\leq i<j\leq\pi(n)}|S_{p_{i}}\cap S_{p_{j}}|-\dots+(-1)^{\pi(n)}\left|\bigcap\limits_{1\leq i\leq\pi(n)}S_{p_{i}}\right|.\\ &=C(n,2)-\sum_{1\leq i\leq\pi(n)}C\left(\left|M_{p_{i}^{\acute{v}_{p_{i}}(n)}}\right|,2\right)+\sum_{1\leq i<j\leq\pi(n)}C\left(\left|M_{p_{i}^{\acute{v}_{p_{i}}(n)}}\cap M_{p_{j}^{\acute{v}_{p_{j}}(n)}}\right|,2\right)\\ &\qquad\qquad\quad-\dots+(-1)^{\pi(n)}C\left(\left|\bigcap\limits_{1\leq i\leq\pi(n)}M_{p_{i}^{\acute{v}_{p_{i}}(n)}}\right|,2\right).\\ &=C(n,2)-\sum_{1\leq i\leq\pi(n)}C\left(\left\lfloor\frac{n}{p_{i}^{\acute{v}_{p_{i}}(n)}}\right\rfloor,2\right)+\hskip-2.84544pt\sum_{1\leq i<j\leq\pi(n)}C\left(\left\lfloor\frac{n}{p_{i}^{\acute{v}_{p_{i}}(n)}p_{j}^{\acute{v}_{p_{j}}(n)}}\right\rfloor,2\right)\\ &\qquad\qquad\quad-\dots+(-1)^{\pi(n)}C\left(\left\lfloor\frac{n}{\prod\limits_{1\leq i\leq\pi(n)}p_{i}^{\acute{v}_{p_{i}}(n)}}\right\rfloor,2\right).\hskip 116.65646pt\square\end{split}$
 4 Degrees of Vertices
+In this section, we proved that there are infinitely many Diophantine graphs which are not prime graphs using a characterization of formula for full degree vertices.
+Moreover, we give a formula for the degree of vertices.
+Necessary and sufficient conditions of equal degree vertices with specific labels and specific reduced labels are investigated.
 4.1 The Degree of a Vertex in $D_{n}$
 Theorem 4.1.
-Suppose $D_{n}$ is the maximal Diophantine graph of order $n$ with a labeling $f$. For each $u\in V(D_{n})$, $p$ is a prime number,
-$\deg(u)=n-1\quad\mbox{if and only if}\quad f(u)\mid n\quad\mbox{{or}}\quad\frac{n}{2}<f(u)=p^{\acute{v}_{p}(n)}<n,$
+For each vertex $u$ in $D_{n}$ and prime number $p$, $\deg(u)=n-1$ if and only if
+$f(u)=p^{\acute{v}_{p}(n)},\ \frac{n}{2}<f(u)<n,\quad\mbox{{or}}\quad f(u)\mid n,$
 where the exclusive or will be typed in bold while the inclusive or is typed as usual.
 Proof.
-Assume that $D_{n}$ is the maximal Diophantine graph of order $n$ with a labeling $f$ and $u\in V(D_{n})$ with $\deg(u)=n-1$. By the way of contradiction, suppose
-$f(u)\nmid n\quad\mbox{and}\quad\left(f(u)\neq p^{\acute{v}_{p}(n)}\quad\mbox{or}\quad\frac{n}{2}\nless f(u)=p^{\acute{v}_{p}(n)}\nless n\right),$
-where $p$ is a prime number. In case of $f(u)=p^{\acute{v}_{p}(n)}$ and $1<f(u)<\frac{n}{2}$, we have
-$(f(u),f(v))=(f(u),2f(u))=\left(p^{\acute{v}_{p}(n)},2p^{\acute{v}_{p}(n)}\right)=p^{\acute{v}_{p}(n)}\nmid n,$
-where $v\in V(D_{n})$ such that $f(v)=2f(n)<n$, which contradicts the full degree of $u$.
-In the other case, when $f(u)\neq p^{\acute{v}_{p}(n)}$ and $f(u)\nmid n$, we have that there exists a prime number $p\mid f(u)$ such that $v_{p}(f(u))>v_{p}(n).$ Then we get
-$v_{p}\left(f(u)\right)\geq v_{p}(n)+1=\acute{v}_{p}(n),$
-so $p^{\acute{v}_{p}(n)}\mid f(u)$. Therefore, one can see that
-$\left(f(u),p^{\acute{v}_{p}(n)}\right)=p^{\acute{v}_{p}(n)}\nmid n,$
-which is a contradiction of the full degree of $u$. Therefore, one can see that
-$f(u)\nmid n\quad\mbox{{or}}\quad f(u)=p^{\acute{v}_{p}(n)},\quad\frac{n}{2}<p^{\acute{v}_{p}(n)}<n.$
-Conversely, in the first case of $f(u)\mid n$, we have for each $v\in V(D_{n})$, $(f(u),f(v))\mid n$.
-In the second case of $f(u)=p^{\acute{v}_{p}(n)}$ and $\frac{n}{2}<f(u)<n$, where $p$ is a prime number, we get any proper multiple of $f(u)$ is bigger than $n$.
+Let $u\in V(D_{n})$ with $\deg(u)=n-1$. By the way of contradiction, suppose
+$f(u)=p^{\acute{v}_{p}(n)},\ \frac{n}{2}<f(u)<n,\quad\mbox{and}\quad f(u)\mid n,$
+or
+$\left(f(u)\neq p^{\acute{v}_{p}(n)}\quad\mbox{or}\quad 1\leq f(u)\leq\frac{n}{2}\right)\quad\mbox{and}\quad f(u)\nmid n,$
+where $p$ is a prime number. The first case is clearly a contradictory statement. The second case has two subcases as follows.
+$\left(f(u)\neq p^{\acute{v}_{p}(n)}\quad\mbox{and}\quad f(u)\nmid n\right)\quad\mbox{or}\quad\left(1<f(u)<\frac{n}{2}\quad\mbox{and}\quad f(u)\nmid n\right),$
+In case of $f(u)\neq p^{\acute{v}_{p}(n)}$ and $f(u)\nmid n$, we get that there exists a prime number $p$ such that $p\mid f(u)$ and $v_{p}(f(u))>v_{p}(n).$ Then, we obtain $v_{p}\left(f(u)\right)\geq\acute{v}_{p}(n),$ so $p^{\acute{v}_{p}(n)}\mid f(u)$.
+Consequently, one can see that $\left(f(u),p^{\acute{v}_{p}(n)}\right)=p^{\acute{v}_{p}(n)}\nmid n,$ which is a contradiction of the full degree of $u$.
+In the other case of $1<f(u)=p^{\acute{v}_{p}(n)}<\frac{n}{2}$, we have $(f(u),2p^{\acute{v}_{p}(n)})\nmid n,$ which contradicts the full degree of $u$.
+Hence, one can see that
+$f(u)=p^{\acute{v}_{p}(n)},\ \frac{n}{2}<p^{\acute{v}_{p}(n)}<n\quad\mbox{{or}}\quad f(u)\mid n.$
+Conversely, in the first case of $f(u)\mid n$, we have that for each $v\in V(D_{n})$, $(f(u),f(v))\mid n$.
+In the second case of $f(u)=p^{\acute{v}_{p}(n)}$ and $\frac{n}{2}<f(u)<n$, where $p$ is a prime number, we get that any proper multiple of $f(u)$ is greater than $n$.
 Since $f(u)=p^{\acute{v}_{p}(n)}$, therefore for all $v\in V(D_{n})$, $(f(u),f(v))\mid p^{v_{p}(n)}$, which means for all $v\in V(D_{n})$, $(f(u),f(v))\mid n$.
-In the both cases, we get $\deg(u)=n-1$.
+In both cases, we get $\deg(u)=n-1$.
 ∎
-Similar results can be deduced in the case of the maximal prime graphs $R_{n}$ for Lemma 3.1 and Theorem 4.1.
-The properties of the maximal Diophantine graphs $D_{n}$ given in Lemma 3.1 and Theorem 4.1 are essential in what follows.
+Similar results can be deduced in the case of a maximal prime graphs $R_{n}$ for Lemma 3.1 and Theorem 4.1.
+In light of Lemma 3.1, Theorem 4.1, and the following theorem 4.2 due to M.
+A.
+Seoud and M.
+Z.
+Youssef, we have the following results.
+Theorem 4.2.
+[11], [13] If $G$ is a simple graph of $n$ vertices which has more than
+$\pi\left(n\right)+\pi\left(\frac{n}{2}\right)+1$
+vertices of degree n1n-1, then $G$ is not prime, where $\pi(n)$ is the number of prime numbers not exceeding $n$.
 Corollary 4.1.
@@
 Proof.
-Let $D_{p^{k}}$ be the maximal Diophantine graphs with $p^{k}$ vertices, where $k>1$ and $p$ is a prime number.
-$D_{p^{k}}$ can not be a prime graph because the number full degree vertices in $R_{p^{k}}$ that is equal to
+Let $p$ be a prime number and an integer $k>1$.
+Maximal Diophantine graphs $D_{p^{k}}$ with $p^{k}$ vertices can not be prime graphs because the number of full degree vertices in $R_{p^{k}}$ that is equal to
 […]
-is less than of the number full degree vertices in $D_{p^{k}}$ which is at least
+and it is less than the number of full degree vertices in $D_{p^{k}}$ which is at least
 […]
@@
 Lemma 4.1.
-Suppose $D_{n}$ is the maximal Diophantine graph of order $n$ with a labeling $f$ and $u\in V(D_{n})$. If $f^{*}(u)=p^{k}$, for some $p$ is a prime number and $k\in\mathbb{N}$, then
+Let $u$ be a vertex in $D_{n}$. If $f^{*}(u)=p^{k}$ for some prime number $p$ and $k\in\mathbb{N}$, then
 $\deg(u)=\left\{\begin{array}[]{ll}n-1,&\hbox{$k=0.$}\\ n-\Big\lfloor\frac{n}{p^{\acute{v}_{p}(n)}}\Big\rfloor,&\hbox{$k>0.$}\end{array}\right$.
 Proof.
-Let $D_{n}$ be the maximal Diophantine graph of order $n$ with a labeling $f$.
-Let $u\in V(D_{n})$ such that $f^{*}(u)=p^{k}$, where $p$ is a prime number and $k\geq 0$.
+Let $f^{*}(u)=p^{k}$, where $u\in V(D_{n})$, $p$ is a prime number and $k\in\mathbb{N}$.
 In the case of $k=0$, we have $f^{*}(u)=1$.
@@
 […]
-Therefore,
-$\deg(u)=n-1-\Big{(}\Big{|}M_{p^{\acute{v}_{p}(n)}}\Big{|}-1\Big{)}=n-\left\lfloor\frac{n}{p^{\acute{v}_{p}(n)}}\right\rfloor.$
-Hence, we have the proof. ∎
-Example 4.1.
-In $D_{20}$, let $f(u)=18$ be the label of a vertex $u\in D_{20}$. Then $f^{*}(u)=9=3^{2}$ and we have $k=2>0$. Therefore, Using Lemma 4.1, we get the following
-$\deg(u)=20-\left\lfloor\frac{20}{3^{\acute{v}_{3}(20)}}\right\rfloor=14.$
-Another way to compute the degree of $u$ with label $18$.
-The neighborhood of $u$ is the set of adjacent vertices of $u$ that has labels belong to the set $\{1,2,4,5,7,8,10,11,13,14,16,17,19,20\}$.
-Therefore, $\deg(u)=14$.
+Therefore, $\deg(u)=n-1-\left(\left|M_{p^{\acute{v}_{p}(n)}}\right|-1\right)=n-\left\lfloor\frac{n}{p^{\acute{v}_{p}(n)}}\right\rfloor.$
 In a very similar way of Lemma 4.1, using the inclusion-exclusion principle and the identity $\sum\limits_{i=0}^{r}(-1)^{r}\binom{r}{i}=0$, one can prove the following theorem.
 Theorem 4.3.
-Suppose $D_{n}$ is the maximal Diophantine graph with order $n$, $f$ is a Diophantine labeling and $u\in V(D_{n})$.
-If $f^{*}(u)=\prod\limits_{i=1}^{r}p_{i}^{k_{i}}$, where $p_{i}$, $i=1,2,\dots,r$ are distinct prime numbers and $k_{i}\geq 0$, $i=1,2,\dots,r$, then
-$\deg(u)=\left\{\begin{array}[]{ll}n-1,&\hbox{$f^{*}(u)=1.$}\\ n-\sum\limits_{1\leq i\leq r}\bigg{\lfloor}\frac{n}{p_{i}^{\acute{v}_{p_{i}}(n)}}\bigg{\rfloor}+\sum\limits_{1\leq i.j\leq r}\bigg{\lfloor}\frac{n}{p_{i}^{\acute{v}_{p_{i}}(n)}p_{j}^{\acute{v}_{p_{j}}(n)}}\bigg{\rfloor}-\dots+(-1)^{r}\Bigg{\lfloor}\frac{n}{\prod\limits_{1\leq i\leq r}p_{i}^{\acute{v}_{p_{i}}(n)}}\Bigg{\rfloor},&\hbox{$f^{*}(u)>1.$}\end{array}\right.$
-Similar results of Theorems 3.1, 4.2 can be obtained in the maximal prime graph $R_{n}$, by putting $v_{p}(n)=0$.
-4.2 Vertices of Equal Degrees $<n-1$
+Let $u$ be a vertex in $D_{n}$. If $f^{*}(u)=\prod\limits_{i=1}^{r}p_{i}^{k_{i}}$, where $p_{i}$, $i=1,2,\dots,r$ are distinct prime numbers and $k_{i}\in\mathbb{N}$, $i=1,2,\dots,r$, then
+$\deg(u)=\left\{\begin{array}[]{ll}n-1,&\\ \hskip 284.52756pt\hbox{if $f^{*}(u)=1.$}\\ n-\sum\limits_{1\leq i\leq r}\left\lfloor\frac{n}{p_{i}^{\acute{v}_{p_{i}}(n)}}\right\rfloor+\sum\limits_{1\leq i.j\leq r}\left\lfloor\frac{n}{p_{i}^{\acute{v}_{p_{i}}(n)}p_{j}^{\acute{v}_{p_{j}}(n)}}\right\rfloor-\dots+(-1)^{r}\left\lfloor\frac{n}{\prod\limits_{1\leq i\leq r}p_{i}^{\acute{v}_{p_{i}}(n)}}\right\rfloor,&\\ \hskip 284.52756pt\hbox{if $f^{*}(u)>1.$}\end{array}\right$.
+Results similar to Theorems 3.1, 4.3 can be obtained in a maximal prime graph $R_{n}$, by putting $v_{p}(n)=0$.
+4.2 Vertices of Equal Degrees <n1<n-1
 Lemma 4.2.
-Suppose $D_{n}$ is the maximal Diophantine graph of order $n$ with a labeling $f$ and $u,v\in V(D_{n})$.
-If $f(u)\mid f(v)$, then $N(u)\supseteq N(v)$, where $N(s)$ defines the neighborhood of $s$ as the set of all vertices in $D_{n}$ that join the vertex $s$.
+Let $u,v$ be two vertices in $D_{n}$. If $f(u)\mid f(v)$, then $N(u)-\{v\}\supseteq N(v)-\{u\}$.
 Proof.
@@
 Lemma 4.3.
-Suppose $D_{n}$ is the maximal Diophantine graph of order $n$ with a labeling $f$. Let $u,v\in V(D_{n})$ such that $f(u)\mid f(v)$. If $\deg(u)=\deg(v)$, then $N(u)=N(v)$.
+Let $u,v$ be two vertices in $D_{n}$ such that $f(u)\mid f(v)$. If $\deg(u)=\deg(v)$, then $N(u)-\{v\}=N(v)-\{u\}$.
 Proof.
-Suppose $D_{n}$ is the maximal Diophantine graph of order $n$ with a labeling $f$. Let $u,v\in V(D_{n})$ such that
-$f(u)\mid f(v)\quad\mbox{and}\quad\deg(u)=\deg(v).$
-Let $s\in V(D_{n})-\{u,v\}$ such that $sv\in E(D_{n})$.
+Let $u,v$ be two vertices in $D_{n}$ such that $f(u)\mid f(v)\ \mbox{and}\ \deg(u)=\deg(v).$ Let $s\in V(D_{n})-\{u,v\}$ such that $sv\in E(D_{n})$.
 Then, using Lemma 4.2, $su\in E(D_{n})$.
-Hence, $N(u)\supseteq N(v)$.
+Hence, $N(u)-\{v\}\supseteq N(v)-\{u\}$.
 On the other hand, let $s\in V(D_{n})-\{u,v\}$ such that $su\in E(D_{n})$.
 Suppose by contrary that $sv\notin E(D_{n})$.
-Since $f(u)\mid f(v)$ and using Lemma 4.2, therefore $N(u)\supset N(v)$.
-Then we have $\deg(u)>\deg(v)$, which contradicts the hypothesis of $\deg(u)=\deg(v)$.
+Since $f(u)\mid f(v)$ and using Lemma 4.2, therefore $N(u)-\{v\}\supset N(v)-\{u\}$.
+Then we have $\deg(u)>\deg(v)$, which contradicts the assumption of $\deg(u)=\deg(v)$.
 Therefore, $sv\in E(D_{n})$.
-Consequently, $N(u)\subseteq N(v)$.
-Hence, $N(u)=N(v)$.
+Consequently, $N(u)-\{v\}\subseteq N(v)-\{u\}$.
+Hence, $N(u)-\{v\}=N(v)-\{u\}$.
 ∎
-It is clear that, if $N(u)=N(v)$, then $\deg(u)=\deg(v)$. Thus, under the assumption that $f(u)\mid f(v)$, we have
-$N(u)=N(v)\quad\mbox{ if and only if}\quad\deg(u)=\deg(v).$
-Furthermore, the following example shows that the condition $f(u)\mid f(v)$ is necessary for Lemma 4.3.
+It is clear that if $N(u)-\{v\}=N(v)-\{u\}$, then $\deg(u)=\deg(v)$.
+Thus, under the assumption that $f(u)\mid f(v)$, we have $N(u)-\{v\}=N(v)-\{u\}$ if and only if $\deg(u)=\deg(v).$ Furthermore, the following example shows that the condition $f(u)\mid f(v)$ is necessary for Lemma 4.3.
 Example 4.1.
 In $D_{20}$, let $f(u)=7$ and $f(v)=8$, where $u,v\in V(D_{20})$.
-$N(u)=V(D_{20})-\{u_{1},u\}$, where $u_{1}\in V(D_{20})$ such that $f(u_{1})=14$ and $N(v)=V(D_{20})-\{v_{1},v\}$, where $v_{1}\in V(D_{20})$ such that $f(v_{1})=16$.
-Then we have $\deg(u)=\deg(v)$ while $f(u)\nmid f(v)$, but $N(u)\neq N(v)$.
+Then $N(u)=V(D_{20})-\{u_{1},u\}$, where $u_{1}\in V(D_{20})$ such that $f(u_{1})=14$ and $N(v)=V(D_{20})-\{v_{1},v\}$, where $v_{1}\in V(D_{20})$ such that $f(v_{1})=16$.
+Then we have $\deg(u)=\deg(v)$, while $f(u)\nmid f(v)$, but $N(u)-\{v\}\neq N(v)-\{u\}$.
 Lemma 4.4.
-Suppose $D_{n}$ is the maximal Diophantine graph of order $n$ with a labeling $f$ and $u\in V(D_{n})$. If $deg(u)<n-1$, then there exists a prime number $p$ such that $p^{\acute{v}_{p}(n)}\mid f(u)$.
+Let $u$ be a vertex in $D_{n}$. If deg(u)<n1deg(u)<n-1, then there exists a prime number $p$ such that $p^{\acute{v}_{p}(n)}\mid f(u)$.
 Proof.
-Suppose $D_{n}$ is the maximal Diophantine graph of order $n$ with a labeling $f$ and let $u\in V(D_{n})$ with $deg(u)<n-1$.
+Let $u$ be a vertex in $D_{n}$ with deg(u)<n1deg(u)<n-1.
 Then there exists a vertex $s\in V(D_{n})$ such that $su\notin E(D_{n})$.
@@
 Theorem 4.4.
-Suppose $D_{n}$ is the maximal Diophantine graph of order $n$ with a labeling $f$.
-Let $f(u)=p^{t}$ and $f(v)=q^{k}$ for some $t,k\geq 0$, where $u,v\in V(D_{n})$ with $\deg(u)<n-1$, $\deg(v)<n-1$ and $p,q$ are two prime numbers.
-Then we have
+Let $f(u)=p^{t}$ and $f(v)=q^{k}$ for some $t,k\in\mathbb{Z^{+}}$, where $u,v\in V(D_{n})$ with $\deg(u)<n-1$, $\deg(v)<n-1$ and $p,q$ are prime numbers. Then we have
 $\deg(u)=\deg(v)\quad\mbox{if and only if}\quad\acute{v}_{p}(n)\leq t,\ \acute{v}_{q}(n)\leq k\quad\mbox{and}\quad\left\lfloor\frac{n}{p^{\acute{v}_{p}(n)}}\right\rfloor=\left\lfloor\frac{n}{q^{\acute{v}_{q}(n)}}\right\rfloor.$
 Proof.
-Suppose $D_{n}$ is the maximal Diophantine graph of order $n$ with a labeling $f$.
-Let $f(u)=p^{t}$ and $f(v)=q^{k}$ for some $t,k>0$, where $p,q$ are primes numbers and $u$ $v\in V(D_{n})$ with $deg(u)=deg(v)$, $deg(u)<n-1$, $deg(v)<n-1$.
+Let $f(u)=p^{t}$ and $f(v)=q^{k}$ for some $t,k\in\mathbb{Z^{+}}$, where $p,q$ are prime numbers and $u,v\in V(D_{n})$ with $deg(u)=deg(v)$, deg(u)<n1deg(u)<n-1 and deg(v)<n1deg(v)<n-1.
 Therefore, using Lemma 4.4, we have
 $p^{\acute{v}_{p}(n)}\mid f(u)=p^{t}\quad\mbox{and}\quad q^{\acute{v}_{q}(n)}\mid f(v)=q^{k}.$
-Consequently, $\acute{v}_{p}(n)\leq t$ and $\acute{v}_{q}(n)\leq k$. By using Lemma 4.1, we get
+Consequently, $\acute{v}_{p}(n)\leq t$ and $\acute{v}_{q}(n)\leq k$. Using Lemma 4.1, we get
 $deg(u)=n-\left\lfloor\frac{n}{p^{\acute{v}_{p}(n)}}\right\rfloor\quad\mbox{and}\quad\deg(v)=n-\left\lfloor\frac{n}{q^{\acute{v}_{q}(n)}}\right\rfloor.$
-Hence, we obtain the following equation
-$\left\lfloor\frac{n}{p^{\acute{v}_{p}(n)}}\right\rfloor=\left\lfloor\frac{n}{q^{\acute{v}_{q}(n)}}\right\rfloor.$
-Conversely, let $\left\lfloor\frac{n}{p^{\acute{v}_{p}(n)}}\right\rfloor=\left\lfloor\frac{n}{q^{\acute{v}_{q}(n)}}\right\rfloor$ and $\acute{v}_{p}(n)\leq t$, $\acute{v}_{q}(n)\leq k$.
-Then we get $p^{\acute{v}_{p}(n)}\mid f(u)\ \mbox{and}\ q^{\acute{v}_{q}(n)}\mid f(v).$ Thus, using Lemma 4.1, we have the following two equations
+Hence, we obtain the following equation $\left\lfloor\frac{n}{p^{\acute{v}_{p}(n)}}\right\rfloor=\left\lfloor\frac{n}{q^{\acute{v}_{q}(n)}}\right\rfloor.$
+Conversely, let $\left\lfloor\frac{n}{p^{\acute{v}_{p}(n)}}\right\rfloor=\left\lfloor\frac{n}{q^{\acute{v}_{q}(n)}}\right\rfloor$ and $\acute{v}_{p}(n)\leq t$, $\acute{v}_{q}(n)\leq k$. Then we get
+$p^{\acute{v}_{p}(n)}\mid f(u)\quad\mbox{and}\quad q^{\acute{v}_{q}(n)}\mid f(v).$
+Thus, using Lemma 4.1, we have the following two equations
 $\deg(u)=n-\left\lfloor\frac{n}{p^{\acute{v}_{p}(n)}}\right\rfloor\quad\mbox{and}\quad\deg(v)=n-\left\lfloor\frac{n}{q^{\acute{v}_{q}(n)}}\right\rfloor.$
-Hence, the result follows. ∎
+Hence, the proof follows. ∎
 Corollary 4.2.
-Suppose $D_{n}$ is the maximal Diophantine graph of order $n$ with a labeling $f.$ Let $u,v$ be two vertices with $\deg(u)<n-1$, $\deg(v)<n-1$ such that $f(u)=p^{t}$, $f(v)=p^{k}$ for some $t,k\geq 0$, where $p$ is a prime number.
-Then we have
+Let $f(u)=p^{t}$ and $f(v)=p^{k}$ for some $t,k\in\mathbb{Z^{+}}$, where $u,v\in V(D_{n})$ with $\deg(u)<n-1$, $\deg(v)<n-1$ and $p$ is a prime number. Then we have
 $\deg(u)=\deg(v)\quad\mbox{if and only if}\quad\acute{v}_{p}(n)\leq\min\{t,k\}.$
 Theorem 4.5.
-Suppose $D_{n}$ is the maximal Diophantine graph of order $n$ with a labeling $f.$ Let $u,v$ be two vertices such that $f(v)=tf(u)$ for some $t\geq 1$ and $f(v)$ is not a prime power. Then we have
-$\deg(u)=\deg(v)\quad\mbox{if and only if for every prime number}\ p\mid t,\quad\acute{v}_{p}(n)\leq v_{p}(f(u))\quad\mbox{{or}}\quad\acute{v}_{p}(n)>v_{p}((f(v))).$
+Let $u,v$ be two vertices in $D_{n}$ such that $f(v)=tf(u)$ for some $t\in\mathbb{Z^{+}}$ and $f(v)$ is not a prime power.
+Then we have $\deg(u)=\deg(v)$ if and only if for every prime number $p\mid t$,
+$\acute{v}_{p}(n)\leq v_{p}(f(u))\quad\mbox{{or}}\quad\acute{v}_{p}(n)>v_{p}(f(v)).$
 Proof.
-Suppose $D_{n}$ is the maximal Diophantine graph of order $n$ with a labeling $f$.
-Let $u,v\in V(D_{n})$ such that $f(v)=tf(u)$ for some $t\geq 1$, $f(v)$ is not a prime power.
-Proof by contrapositive, suppose that there exists a prime number $p_{0}\mid t$ such that
-$\acute{v}_{p_{0}}(n)\leq v_{p_{0}}(f(u))\quad\mbox{and}\quad\acute{v}_{p_{0}}(n)>v_{p_{0}}((f(v))),$
+Let $u,v$ be two vertices in $D_{n}$ such that $f(v)=tf(u)$ for some $t\in\mathbb{Z^{+}}$ and $f(v)$ is not a prime power.
+Proof by contraposition, suppose that there exists a prime number $p_{0}\mid t$ such that
+$\acute{v}_{p_{0}}(n)\leq v_{p_{0}}(f(u))\quad\mbox{and}\quad\acute{v}_{p_{0}}(n)>v_{p_{0}}(f(v)),$
 or
 $\acute{v}_{p_{0}}(n)>v_{p_{0}}(f(u))\quad\mbox{and}\quad\acute{v}_{p_{0}}(n)\leq v_{p_{0}}(f(v)).$
-The first case contradicts the assumptions $f(u)\mid f(v)$. Therefore, let us consider the second case, which is
+The first case contradicts assumptions $f(u)\mid f(v)$. Therefore, let us consider the second case, which is
 $v_{p_{0}}(f(u))<\acute{v}_{p_{0}}(n)\leq v_{p_{0}}(f(v)).$ (4)
 Let a vertex $s\in V(D_{n})$ such that $f(s)=p_{0}^{\acute{v}_{p_{0}}(n)}$.
-Therfore, using equation (4), $p_{0}^{\acute{v}_{p_{0}}(n)}\nmid p_{0}^{\acute{v}_{p_{0}}(f(u))}$.
-Since $p_{0}^{\acute{v}_{p_{0}}(f(u))}\mid f(u)$, we get $f(s)=p_{0}^{\acute{v}_{p_{0}}(n)}\nmid f(u)$.
+Then, using equation (4), we have $p_{0}^{\acute{v}_{p_{0}}(n)}\nmid p_{0}^{\acute{v}_{p_{0}}(f(u))}.$ Since $p_{0}^{\acute{v}_{p_{0}}(f(u))}\mid f(u)$, we get $f(s)=p_{0}^{\acute{v}_{p_{0}}(n)}\nmid f(u)$.
 Consequently, $s\neq u$.
-Also, $s\neq v$ for $f(v)$ is not a prime power.
+In addition, $s\neq v$, since $f(v)$ is not a prime power.
 Moreover, using equation (4), we get
 $(f(u),f(s))\mid n\quad\mbox{while}\quad(f(v),f(s))\nmid n.$
-Thus,
-$su\in E(D_{n})\quad\mbox{while}\quad sv\notin E(D_{n}).$
-Since $f(u)\mid f(v)$ and using Lemma 4.2, we have $N(u)\supset N(v)$. Hence, $\deg(u)\neq\deg(v)$.
-Conversely, suppose that, by contrapositive, $deg(u)\neq deg(v)$. Then, using Lemma 4.2, $N(u)\supset N(v)$. So, there exists a vertex $s\in V(D_{n})-\{u,v\}$ such that
-$s\in N(u)\quad\mbox{and}\quad s\notin N(v).$
-Consequently,
+Thus, $su\in E(D_{n})\ \mbox{while}\ sv\notin E(D_{n}).$ Since $f(u)\mid f(v)$ and using Lemma 4.2, we have $N(u)-\{v\}\supset N(v)-\{u\}$. Hence, $\deg(u)\neq\deg(v)$.
+Conversely, suppose that, by contrapositive, $deg(u)\neq deg(v)$.
+Then, using Lemma 4.2, $N(u)-\{v\}\supset N(v)-\{u\}$.
+So, there exists a vertex $s\in V(D_{n})-\{u,v\}$ such that $s\in N(u)-\{v\}\ \mbox{and}\ s\notin N(v)-\{u\}.$ Consequently,
 $(f(u),f(s))\mid n\quad\mbox{and}\quad(f(v),f(s))\nmid n.$ (5)
 Then, using Lemma 3.1, there exists a prime number $p_{0}$ such that
-$f(v),f(s)\in M_{p_{0}^{\acute{v}_{p_{0}}(n)}}\quad\mbox{and}\quad p_{0}^{\acute{v}_{p_{0}}(n)}\nmid f(u).$ (6)
-Let $f(s)=p_{0}^{\acute{v}_{p_{0}}(n)}$. Then, using equation 6, we have $s\neq u$. Since $f(v)=tf(u)$, we get $p_{0}\mid t$ and $s\neq v$. Therefore, using equation 5, we have the following
+$p_{0}^{\acute{v}_{p_{0}}(n)}\nmid f(u)\quad\mbox{and}\quad f(v),f(s)\in M_{p_{0}^{\acute{v}_{p_{0}}(n)}}.$ (6)
+Let $f(s)=p_{0}^{\acute{v}_{p_{0}}(n)}$. Then, using equation (6), we have $s\neq u$. Since $f(v)=tf(u)$, we get $p_{0}\mid t$ and $s\neq v$. Therefore, using equation (5), we have
 $v_{p_{0}}\left((f(u),f(s))\right)\leq v_{p_{0}}(n)\quad\mbox{and}\quad v_{p_{0}}\left((f(v),f(s))\right)>v_{p_{0}}(n).$ (7)
-Hence, using equation 7, we get $v_{p_{0}}(f(u))<\acute{v}_{p_{0}}(n)$ and $v_{p_{0}}(f(v))\geq\acute{v}_{p_{0}}(n)$ for some prime number $p_{0}\mid t$. ∎
+Hence, using equation (7), we get
+$v_{p_{0}}(f(u))<\acute{v}_{p_{0}}(n)\quad\mbox{and}\quad v_{p_{0}}(f(v))\geq\acute{v}_{p_{0}}(n)$
+for some prime number $p_{0}\mid t$. ∎
 4.3 More Necessary and Sufficient Conditions for Equality of Degrees
-The following theorem is a generalization of one direction of Theorems 4.3.
+The following theorem is a generalization of one direction of Theorem 4.4.
 Theorem 4.6.
-Suppose $D_{n}$ is the maximal Diophantine graph of order $n$ with a labeling $f$ and $u,v\in V(D_{n})$ such that $\deg(u)<n-1$, $\deg(v)<n-1$.
-If $\omega\big{(}f^{*}(u)\big{)}=\omega\big{(}f^{*}(v)\big{)}$ and for every prime number $p\mid f^{*}(u)$, there exits a prime number $q\mid f^{*}(v)$ such that $\left\lfloor\frac{n}{p^{\acute{v}_{p}(n)}}\right\rfloor=\left\lfloor\frac{n}{q^{\acute{v}_{q}(n)}}\right\rfloor$, then $\deg(u)=\deg(v)$, where $w(n)$ is the number of prime factors of $n$.
+Let $u,v$ be two vertices in $D_{n}$ with $\deg(u)<n-1$ and $\deg(v)<n-1$.
+If $\omega\big(f^{*}(u)\big)=\omega\big(f^{*}(v)\big)$ and for every prime number $p\mid f^{*}(u)$, there exists a prime number $q\mid f^{*}(v)$ such that $\left\lfloor\frac{n}{p^{\acute{v}_{p}(n)}}\right\rfloor=\left\lfloor\frac{n}{q^{\acute{v}_{q}(n)}}\right\rfloor$, then $\deg(u)=\deg(v)$, where $w(n)$ is the number of distinct prime factors of $n$.
 Proof.
-Suppose $D_{n}$ is the maximal Diophantine graph of order $n$ with a labeling $f$ and $u,v\in V(D_{n})$ such that
-$\deg(u)<n-1,\quad\deg(v)<n-1\quad\mbox{and}\quad\omega\big{(}f^{*}(u)\big{)}=\omega\big{(}f^{*}(v)\big{)}.$
-Therefore, $\omega\big{(}f^{*}(u)\big{)}=\omega\big{(}f^{*}(v)\big{)}>0.$ Hence, there is a prime number $p$ such that $p\mid f^{*}(u)$.
-Consequently, our hypothesis implies that there is a prime number $q\mid f^{*}(v)$ such that
+Let $u,v$ be two vertices in $D_{n}$ such that $\deg(u)<n-1$, $\deg(v)<n-1$ and
+$\omega\big(f^{*}(u)\big)=\omega\big(f^{*}(v)\big).$ (8)
+Then, using Lemma 4.4, there exists a prime number $p$ such that $p^{\acute{v}_{p}(n)}\mid f(u).$ Therefore, using Lemma 2.1, part i., there exists a prime number $p$ such that $p\mid f^{*}(u)$.
+Consequently, It follows from our assumption that there exists a prime number $q$ such that $q\mid f^{*}(v)$ and
 $\left\lfloor\frac{n}{p^{\acute{v}_{p}(n)}}\right\rfloor=\left\lfloor\frac{n}{q^{\acute{v}_{q}(n)}}\right\rfloor.$ (9)
-Therefore, using Lemmas 2.2 part $(i)$, $p^{\acute{v}_{p}(n)}\mid f(u)$ and $q^{\acute{v}_{q}(n)}\mid f(v).$ Then, using the fundamental theorem of arithmetic, there exist $\grave{u},\grave{v}\in V(D_{n})$ such that
-$f(\grave{u}):=\prod\limits_{p\mid f^{*}(u)}p^{\acute{v}_{p}(n)}\ \Big{|}\ f(u)\quad\mbox{and}\quad f(\grave{v}):=\prod\limits_{q\mid f^{*}(v)}q^{\acute{v}_{q}(n)}\ \Big{|}\ f(v).$
-By definition of $f(\grave{u})$ and $f(\grave{v})$, we have $f^{*}(u),f^{*}(\grave{u})$ have the same prime factors and $f^{*}(v),f^{*}(\grave{v})$ have the same prime factors.
-Since $\omega\big{(}f^{*}(u)\big{)}=\omega\big{(}f^{*}(v)\big{)},$ therefore
-$\omega\big{(}f^{*}(\grave{u})\big{)}=\omega\big{(}f^{*}(u)\big{)}=\omega\big{(}f^{*}(v)\big{)}=\omega\big{(}f^{*}(\grave{v})\big{)}.$ (9)
-Moreover, for every prime number $p$, we have
-$p^{\acute{v}_{p}(n)}\mid f(u)\ \mbox{if and only if}\ p^{\acute{v}_{p}(n)}\mid f(\grave{u}).$
+Moreover, $\omega\big(f^{*}(u)\big)=\omega\big(f^{*}(v)\big)>0.$ Since $p^{\acute{v}_{p}(n)}\mid f(u)$ and $q^{\acute{v}_{q}(n)}\mid f(v)$ and using Theorem 1.1, therefore there exist $\grave{u},\grave{v}\in V(D_{n})$ such that
+$f(\grave{u}):=\prod\limits_{p\mid f^{*}(u)}p^{\acute{v}_{p}(n)}\ \bigg|\ f(u)\quad\mbox{and}\quad f(\grave{v}):=\prod\limits_{q\mid f^{*}(v)}q^{\acute{v}_{q}(n)}\ \bigg|\ f(v).$
+[…]
+$).\omega\big(f^{*}(\grave{u})\big)=\omega\big(f^{*}(u)\big)=\omega\big(f^{*}(v)\big)=\omega\big(f^{*}(\grave{v}$$\big)$. (10)
+Furthermore, for every prime number $p$, we have
+[…]
 Therefore,
@@
 Hence,
-$\deg(u)=\deg(\grave{u}).$ (10)
+[…]
 Similarly, we get
-$\deg(v)=\deg(\grave{v}).$ (11)
-Let $d=(f(\grave{u}),f(\grave{v}))$. Thus, Lemmas 2.2 part $(ii)$ implies that
-$\frac{d}{(d,n)}\ \bigg{|}\ f^{*}(\grave{u}),f^{*}(\grave{v}),$ (12)
-and hence
-$\omega\left(\frac{d}{(d,n)}\right)\leq\omega\big{(}f^{*}(\grave{u})\big{)}=\omega\big{(}f^{*}(\grave{v})\big{)}.$
-Let us consider the following cases.
-Case 1: Let $\omega\big{(}f^{*}(\grave{u})\big{)}=\omega\big{(}f^{*}(\grave{v})\big{)}=\omega\left(\frac{d}{(d,n)}\right).$ Then, using equation (12), $\frac{d}{(d,n)}$, $f^{*}(\grave{u})$ and $f^{*}(\grave{v})$ have the same prime factors.
-A similar argument used to prove equation (10) implies that
-$\deg(\grave{u})=\deg(\grave{v}).$
-Hence, using equations (10) and (11), we get
-$\deg(u)=\deg(v).$
-Case 2: Let $\omega\big{(}f^{*}(\grave{u})\big{)}=\omega\big{(}f^{*}(\grave{v})\big{)}=\omega\left(\frac{d}{(d,n)}\right)+1=r.$ Then we get $f^{*}(\grave{u})$ and $f^{*}(\grave{v})$ differ in exactly one prime factor.
-Therefore, using the fundamental theorem of arithmetic, we obtain the following two equations
+[…]
+Let $)d=(f(\grave{u}),f(\grave{v}$)). Thus, Lemmas 2.1, part ii.ii., implies that
+[…]
+[…]
+[…]
+[…]
 $f(\grave{u})=\prod\limits_{i=1}^{r-1}h_{i}^{\acute{v}_{h_{i}}(n)}p^{\acute{v}_{p}(n)}\quad\mbox{and}\quad f(\grave{v})=\prod\limits_{i=1}^{r-1}h_{i}^{\acute{v}_{h_{i}}(n)}q^{\acute{v}_{q}(n)},$
 where $h_{i},p$ and $q$ are distinct prime numbers such that $h_{i}\mid d,p\nmid d$ and $q\nmid d$ for all $i=1,2,\dots,r-1$. Then, using Theorem 4.3, we have the following two equations
-$\begin{split}\deg(\grave{u})=n-\sum_{1\leq i\leq r-1}\left\lfloor\frac{n}{h_{i}^{\acute{v}_{h_{i}}(n)}}\right\rfloor&-\left\lfloor\frac{n}{p^{\acute{v}_{p}(n)}}\right\rfloor+\sum_{1\leq i<j\leq r-1}\left\lfloor\frac{n}{h_{i}^{\acute{v}_{h_{i}}(n)}h_{j}^{\acute{v}_{h_{j}}(n)}}\right\rfloor+\sum_{1\leq i\leq r-1}\left\lfloor\frac{n}{h_{i}^{\acute{v}_{h_{i}}(n)}p^{\acute{v}_{p}(n)}}\right\rfloor-\\ &-\dots+(-1)^{r}\left\lfloor\frac{n}{\prod\limits_{1\leq i\leq r-1}h_{i}^{\acute{v}_{h_{i}}(n)}p^{\acute{v}_{p}(n)}}\right\rfloor,\end{split}$ (13)
+$\begin{split}\deg(\grave{u})=n&-\sum_{1\leq i\leq r-1}\left\lfloor\frac{n}{h_{i}^{\acute{v}_{h_{i}}(n)}}\right\rfloor-\left\lfloor\frac{n}{p^{\acute{v}_{p}(n)}}\right\rfloor+\sum_{1\leq i<j\leq r-1}\left\lfloor\frac{n}{h_{i}^{\acute{v}_{h_{i}}(n)}h_{j}^{\acute{v}_{h_{j}}(n)}}\right\rfloor\\ &+\sum_{1\leq i\leq r-1}\left\lfloor\frac{n}{h_{i}^{\acute{v}_{h_{i}}(n)}p^{\acute{v}_{p}(n)}}\right\rfloor-\dots+(-1)^{r}\left\lfloor\frac{n}{\prod\limits_{1\leq i\leq r-1}h_{i}^{\acute{v}_{h_{i}}(n)}p^{\acute{v}_{p}(n)}}\right\rfloor,\end{split}$ (14)
 and
-$\begin{split}\deg(\grave{v})=n-\sum_{1\leq i\leq r-1}\left\lfloor\frac{n}{h_{i}^{\acute{v}_{h_{i}}(n)}}\right\rfloor&-\left\lfloor\frac{n}{q^{\acute{v}_{q}(n)}}\right\rfloor+\sum_{1\leq i<j\leq r-1}\left\lfloor\frac{n}{h_{i}^{\acute{v}_{h_{i}}(n)}h_{j}^{\acute{v}_{h_{j}}(n)}}\right\rfloor+\sum_{1\leq i\leq r-1}\left\lfloor\frac{n}{h_{i}^{\acute{v}_{h_{i}}(n)}q^{\acute{v}_{q}(n)}}\right\rfloor-\\ &-\dots+(-1)^{r}\left\lfloor\frac{n}{\prod\limits_{1\leq i\leq r-1}h_{i}^{\acute{v}_{h_{i}}(n)}q^{\acute{v}_{q}(n)}}\right\rfloor.\end{split}$ (14)
-Using Lemma 2.6 and equations (8), (13) and (14), we get
-$\deg(\grave{u})=\deg(\grave{v}),$
-and hence
-$\deg(u)=\deg(v).$
-Case 3: Let $\omega\big{(}f^{*}(\grave{u})\big{)}=\omega\big{(}f^{*}(\grave{v})\big{)}=\omega\left(\frac{d}{(d,n)}\right)+2=r.$ Then $f^{*}(\grave{u})$ and $f^{*}(\grave{v})$ differ in exactly two prime factors.
-Therefore, using the fundamental theorem of arithmetic, we have the following two equations
+$\begin{split}\deg(\grave{v})=n&-\sum_{1\leq i\leq r-1}\left\lfloor\frac{n}{h_{i}^{\acute{v}_{h_{i}}(n)}}\right\rfloor-\left\lfloor\frac{n}{q^{\acute{v}_{q}(n)}}\right\rfloor+\sum_{1\leq i<j\leq r-1}\left\lfloor\frac{n}{h_{i}^{\acute{v}_{h_{i}}(n)}h_{j}^{\acute{v}_{h_{j}}(n)}}\right\rfloor\\ &+\sum_{1\leq i\leq r-1}\left\lfloor\frac{n}{h_{i}^{\acute{v}_{h_{i}}(n)}q^{\acute{v}_{q}(n)}}\right\rfloor-\dots+(-1)^{r}\left\lfloor\frac{n}{\prod\limits_{1\leq i\leq r-1}h_{i}^{\acute{v}_{h_{i}}(n)}q^{\acute{v}_{q}(n)}}\right\rfloor.\end{split}$ (15)
+[…]
+[…]
 $f(\grave{u})=\prod\limits_{i=1}^{r-2}h_{i}^{\acute{v}_{h_{i}}(n)}p_{1}^{\acute{v}_{p_{1}}(n)}p_{2}^{\acute{v}_{p_{2}}(n)}\quad\mbox{and}\quad f(\grave{v})=\prod\limits_{i=1}^{r-2}h_{i}^{\acute{v}_{h_{i}}(n)}q_{1}^{\acute{v}_{q_{1}}(n)}q_{2}^{\acute{v}_{q_{1}}(n)},$
@@
 $\left\lfloor\frac{n}{p_{1}^{\acute{v}_{p_{1}}(n)}}\right\rfloor=\left\lfloor\frac{n}{q_{1}^{\acute{v}_{q_{1}}(n)}}\right\rfloor\quad\mbox{and}\quad\left\lfloor\frac{n}{p_{2}^{\acute{v}_{p_{2}}(n)}}\right\rfloor=\left\lfloor\frac{n}{q_{2}^{\acute{v}_{q_{2}}(n)}}\right\rfloor.$
-Define
-$a:=p_{1}^{\acute{v}_{p_{1}}(n)}p_{2}^{\acute{v}_{p_{2}}(n)}\quad\mbox{and}\quad b:=q_{1}^{\acute{v}_{q_{1}}(n)}q_{2}^{\acute{v}_{q_{1}}(n)}.$
-Since
-$\frac{a}{(a,n)}=p_{1}p_{2}\quad\mbox{and}\quad\frac{b}{(b,n)}=q_{1}q_{2},$
-therefore Lemma 2.5 implies that $a>n\ \mbox{or}\ b>n.$ Hence,
-$f(\grave{u})\geq a>n\quad\mbox{or}\quad f(\grave{v})\geq b>n,$
-which is absurd.
-The cases $\omega\big{(}f^{*}(\grave{u})\big{)}=\omega\big{(}f^{*}(\grave{v})\big{)}=\omega\left(\frac{d}{(d,n)}\right)+i$, for every $i\geq 3$ can be treated in a very similar way of case $3$ which are absurd.
-Hence the result follows.
+Define $a:=p_{1}^{\acute{v}_{p_{1}}(n)}p_{2}^{\acute{v}_{p_{2}}(n)}\ \mbox{and}\ b:=q_{1}^{\acute{v}_{q_{1}}(n)}q_{2}^{\acute{v}_{q_{1}}(n)}.$ Since $\frac{a}{(a,n)}=p_{1}p_{2}\ \mbox{and}\ \frac{b}{(b,n)}=q_{1}q_{2},$ therefore Lemma 2.4 implies that $a>n\ \mbox{or}\ b>n.$ Hence, $f(\grave{u})\geq a>n\ \mbox{or}\ f(\grave{v})\geq b>n,$ which is absurd.
+The cases $\omega\big(f^{*}(\grave{u})\big)=\omega\big(f^{*}(\grave{v})\big)=\omega\left(\frac{d}{(d,n)}\right)+i$, for every $i>3$ can be treated in a very similar way of case $3$ which are absurd.
+Hence the proof follows.
 ∎
 Example 4.2.
-In Example 4.2, let $f(u)=14$ and $f(v)=16$, where $u,v\in V(D_{20})$, so $f^{*}(u)=7$ and $f^{*}(v)=8$.
+In $D_{20}$, let $f(u)=14$ and $f(v)=16$, where $u,v\in V(D_{20})$, so $f^{*}(u)=7$ and $f^{*}(v)=8$.
 Then, applying Theorem 4.6, we have $\omega\left(7\right)=\omega\left(8\right)=1$ and $\left\lfloor\frac{20}{7}\right\rfloor=\left\lfloor\frac{20}{8}\right\rfloor=2.$ Hence, $\deg(u)=\deg(v)=18$, (by applying Lemma 4.1).
-The converse of Theorem 4.5 does not hold.
-For instance, let $u,v\in V(D_{23})$ such that $f(u)=10$, $f(v)=14$.
-By Theorem 4.2, we have $deg(u)=deg(v)=10$.
-However, it can be verified that $\left\lfloor\frac{23}{5}\right\rfloor\neq\left\lfloor\frac{23}{7}\right\rfloor$, which demonstrates that the equality of degrees does not imply the equality of values of the floor function.
+The converse of Theorem 4.6 is not true.
+For instance, in $D_{23}$, let $f(u)=10$, $f(v)=14$, where $u,v\in V(D_{23})$.
+Then, by Theorem 4.3, we have $deg(u)=13=deg(v)$.
+However, it can be verified that $\left\lfloor\frac{23}{5}\right\rfloor\neq\left\lfloor\frac{23}{7}\right\rfloor$.
 Corollary 4.3.
-Suppose $D_{n}$ is the maximal Diophantine graph of order $n$ with a labeling $f$.
-Let $u,v\in V(D_{n})$ such that $f^{*}(u)>1$, $f^{*}(v)>1$.
-If $f^{*}(u),f^{*}(v)$ have the same prime factors, then $\deg(u)=\deg(v)$.
+Let $u,v$ be two vertices in $D_{n}$ with $f^{*}(u)>1$, $f^{*}(v)>1$. If $f^{*}(u),f^{*}(v)$ have the same prime factors, then $\deg(u)=\deg(v)$.
 Corollary 4.4.
-Suppose $D_{n}$ is the maximal Diophantine graph of order $n$ with a labeling $f$.
-Let $u,v\in V(D_{n})$ such that $f(v)=tf(u)$ for some $t\geq 1$.
-If $t\mid n$ and $(t,f(u))=1$, then $\deg(u)=\deg(v)$.
+Let $u,v$ be two vertices in $D_{n}$ such that $f(v)=tf(u)$ for some $t\in\mathbb{Z^{+}}$. If $(t,f(u))=1$ and $t\mid n$, then $\deg(u)=\deg(v)$.
 Proof.
-Suppose $D_{n}$ is the maximal Diophantine graph of order $n$ with a labeling $f$.
-Let $u,v\in V(D_{n})$ such that $f(v)=tf(u)$ for some $t\geq 1$, $t\mid n$ and $(t,f(u))=1$.
-The reduced labels $f^{*}(v),f^{*}(u)$ satisfy the following equalities.
+Let $u,v$ be two vertices in $D_{n}$ such that $f(v)=tf(u)$ for some $t\in\mathbb{Z^{+}}$, $(t,f(u))=1$ and $t\mid n$. The reduced labels $f^{*}(v),f^{*}(u)$ satisfy the following equalities.
 $f^{*}(v)=\frac{f(v)}{(f(v),n)}=\frac{tf(u)}{(tf(u),n)}=\frac{tf(u)}{t(f(u),n)}=\frac{f(u)}{(f(u),n)}=f^{*}(u).$
-Therefore, the two reduced labels $f^{*}(u),f^{*}(v)$ have the same primes factors or each of them is equal to 1. Hence, using Corollary 4.3 and Theorem 4.1, the result follows. ∎
-The following two examples shows the converse of Corollary 4.3 is not true,
+Therefore, $f^{*}(u),f^{*}(v)$ have the same prime factors or $f^{*}(u)=f^{*}(v)=1$. Hence, using Corollary 4.3 and Theorem 4.1, the proof follows. ∎
+The following two examples show that the converse of Corollary 4.3 is not true.
 Example 4.3.
-Suppose $D_{n}$ is the maximal Diophantine graph of order $n$ with a labeling $f$.
-Let $u,v\in V(D_{n})$ such that $f(u)\neq f(v)$, $f(u)\mid f(v)$, and $f(v)=p^{\acute{v}_{p}(n)}$, $\frac{n}{2}<f(v)<n$, where $p$ is a prime number, using Theorem 4.1, we have
-$\deg(u)=n-1=\deg(v).$
-However, $f^{*}(u)=\frac{f(u)}{(f(u),n)}=1$ and $f^{*}(v)=\frac{p^{\acute{v}_{p}(n)}}{\left(p^{\acute{v}_{p}(n)},n\right)}=p$ have not the same prime factors.
+Let $u,v$ be two different vertices in $D_{n}$ such that $f(u)\mid f(v)$ and $f(v)=p^{\acute{v}_{p}(n)}$, $\frac{n}{2}<f(v)<n$, where $p$ is a prime number.
+Then, using Theorem 4.1, we have $\deg(u)=n-1=\deg(v).$ However, $f^{*}(u)=\frac{f(u)}{(f(u),n)}=1$ and $f^{*}(v)=\frac{p^{\acute{v}_{p}(n)}}{\left(p^{\acute{v}_{p}(n)},n\right)}=p$ have not the same prime factors.
 Example 4.4.
-Suppose $D_{n}$ is the maximal Diophantine graph of even order $n$ with a labeling $f$.
-Let $u,v\in V(D_{n})$ such that $f(u)=2p^{\acute{v}_{p}(n)}$ and $f(v)=2q^{\acute{v}_{q}(n)}$, $\frac{n}{3}<p^{\acute{v}_{p}(n)},q^{\acute{v}_{q}(n)}<\frac{n}{2}$, where $p,q$ are distinguish prime numbers, this means $f(u)\nmid f(v)$, using Lemma 4.1 and Lemmas 2.2 part $(i)$, we have
+Let $u,v$ be two vertices in $D_{n}$ such that $f(u)=2p^{\acute{v}_{p}(n)}$ and $f(v)=2q^{\acute{v}_{q}(n)}$, $\frac{n}{3}<p^{\acute{v}_{p}(n)},q^{\acute{v}_{q}(n)}<\frac{n}{2}$, where $p,q$ are distinguish prime numbers, this means $f(u)\nmid f(v)$.
+Then, using Lemma 4.1 and Lemmas 2.1, part $i.$, we have
 $\deg(u)=n-\left\lfloor\frac{n}{p^{\acute{v}_{p}(n)}}\right\rfloor=n-2=n-\left\lfloor\frac{n}{q^{\acute{v}_{q}(n)}}\right\rfloor=\deg(v).$
@@
 Theorem 4.7.
-Suppose $D_{n}$ is the maximal Diophantine graph of order $n$ with a labeling $f$.
-Let $u,v\in V(D_{n})$ such that $f(u)\mid f(v)$, $f(v)$ is not a prime power number and $f^{*}(u)>1$.
-If $\deg(u)=\deg(v)$, then $f^{*}(u),f^{*}(v)$ have the same prime factors.
+Let $u,v$ be two vertices in $D_{n}$ such that $f(u)\mid f(v)$, $f(v)$ is not a prime power and $f^{*}(u)>1$. If $\deg(u)=\deg(v)$, then $f^{*}(u),f^{*}(v)$ have the same prime factors.
 Proof.
-Suppose $D_{n}$ is the maximal Diophantine graph of order $n$ with a labeling $f$.
-Let $u,v\in V(D_{n})$ such that $f(u)\mid f(v)$, $f(v)$ is not a prime power number, $\deg(u)=\deg(v)$ and $f^{*}(u)>1$.
-Then, using Lemmas 2.2 part $(ii)$, we have $f^{*}(u)\mid f^{*}(v)$.
+Let $u,v$ be two vertices in $D_{n}$ such that $f(u)\mid f(v)$, $f(v)$ is not a prime power, $\deg(u)=\deg(v)$ and $f^{*}(u)>1$.
+Then, using Lemmas 2.1, part ii.ii., we have $f^{*}(u)\mid f^{*}(v)$.
 Consequently, $f^{*}(v)>1$ and also any prime factor of $f^{*}(u)$ is a prime factor of $f^{*}(v)$.
 Let there exists a prime number $p$ such that $p\mid f^{*}(v)$.
-Then, using Lemmas 2.2 part $(i)$, $p^{\acute{v}_{p}(n)}\mid f(v)$, i.e., $f(v)\in M_{p^{\acute{v}_{p}(n)}}$.
-Let a vertex $s\in V(D_{n})$ such that $f(s)=p^{\acute{v}_{p}(n)}\in M_{p^{\acute{v}_{p}(n)}}.$ Since $f(v)$ is not a prime power number, we get $s\neq v$.
+Then, using Lemmas 2.1, part $i.$, $p^{\acute{v}_{p}(n)}\mid f(v)$, i.e., $f(v)\in M_{p^{\acute{v}_{p}(n)}}$.
+Let a vertex $s\in V(D_{n})$ such that $f(s)=p^{\acute{v}_{p}(n)}\in M_{p^{\acute{v}_{p}(n)}}.$ Since $f(v)$ is divisible by at least two distinct prime numbers, we get $s\neq v$.
 Using Lemma 3.1, we have $(f(v),f(s))\nmid n$.
-Therefore, $s\notin N(v)$.
-Consequently, using Lemma 4.3, we obtain $s\notin N(u)$.
-Then we get $\left(f(u),p^{\acute{v}_{p}(n)}\right)\nmid n$.
-Thus, $p^{\acute{v}_{p}(n)}\mid f(u)$.
-Consequently, using Lemmas 2.2 part $(i)$, $p\mid f^{*}(u)$.
+Therefore, $s\notin N(v)-\{u\}$.
+Consequently, using Lemma 4.3, we obtain $s\notin N(u)-\{v\}$.
+Then we get $\left(f(u),f(s)\right)\nmid n$.
+Thus, $f(s)=p^{\acute{v}_{p}(n)}\mid f(u)$.
+Consequently, using Lemmas 2.1, part $i.$, $p\mid f^{*}(u)$.
 Therefore, for every prime number $p$,
 $p\mid f^{*}(u)\quad\mbox{if and only if}\quad p\mid f^{*}(v).$
-Hence, the two reduced labels $f^{*}(u),f^{*}(v)$ have the same prime factors. ∎
-Future Work
-Give two efficient algorithms that one determines the maximal Diophantine graphs and anther determines whether a given graph is Diophantine or not.
-Furthermore, exploring different variants of Diophantine labeling can further enrich the theory and uncover new connections between graph theory and number theory.
-Declaration
-The authors contributed equally to the manuscripts and approved the submitted scripts.
-We declare that we have no conflicts of interests.
-We confirm that this work is original and has not been published elsewhere, nor is it currently under consideration for publication elsewhere.
+Hence, the proof follows. ∎
+5 Conclusions
+This manuscript introduces the concept of Diophantine labelling and establishes a connection between Diophantine graphs and prime graphs.
+We study further define maximal Diophantine graphs and derives explicit formulas for their edge count and vertex degrees.
+Additionally, we provide necessary and sufficient conditions for the equality of vertex degrees.
+We also utilize number-theoretic concepts, such as p-adic valuation, to analyze vertex labels and the properties of Diophantine graphs.
+Although Diophantine labelling is still primarily theoretical, we think that Diophantine labelling sits at an intersection of graph theory, number theory and combinatorics.
+Furthermore, exploring different variants of Diophantine labelling can further enrich the theory and uncover new connections between these mathematical domains.
+As research advances, we expect that Diophantine labelling has more concrete applications in the future.
+Moreover, the algorithmic approach is promising in many directions, including searching for an efficient algorithm for Diophantine graph and finding maximal Diophantine graphs.
+Acknowledgment
+The author would like to extend their gratitude to the referees for their valuable feedback, insightful comments, and constructive suggestions, which significantly contributed to improving the quality and clarity of this work.